### v1

# Existence proofs of traveling wave solutions on an infinite strip for the suspension bridge equation and proof of orbital stability

Abstract
[…]
Key words. Traveling waves, PDEs on unbounded domains, Stability analysis, Fourier analysis
1 Introduction
The objective of this paper is to prove the existence of traveling wave solutions of the two-dimensional suspension bridge equation on an infinite strip $\Omega\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}{\mathbb{R}}\times(-d_{2},d_{2})$ for $d_{2}>0$. We let $v=v(t,X_{1},X_{2}):{\mathbb{R}}^{+}\times\Omega\to{\mathbb{R}}$ represent the deflection in the downward direction of the surface of the bridge and we let $\Delta^{2}=(\partial_{X_{1}}^{2}+\partial_{X_{2}}^{2})^{2}$. One of the systems used for modeling suspension bridges is (e.g. [8], [18])
$\partial_{t}^{2}v=-\Delta^{2}v-e^{v}+1.$ (1)
Summaries of the historical development of different mathematical models for suspension bridges, both simplistic and advanced, are available in [11] and [13].
In the dynamics of the suspension bridge equation, traveling waves can be observed [2]. In order to look for such solutions, we introduce the parameter $c\in\mathbb{R}$ to denote the wave speed. We consider waves in the $X_{1}$-direction. Hence, we choose the traveling wave ansatz $v(t,X_{1},X_{2})=u(X_{1}-ct,X_{2})=u(x_{1},x_{2})$ such that the function $u:\Omega\to{\mathbb{R}}$ satisfies
$\Delta^{2}u+c^{2}\partial_{x_{1}}^{2}u+e^{u}-1=0,$ (2)
with $u(x)\to 0$ as $|x_{1}|\to\infty$. Moreover, Neumann boundary conditions are imposed in the $x_{2}$-direction:
$\partial_{x_{2}}u(\cdot,\pm d_{2})=\partial_{x_{2}}^{3}u(\cdot,\pm d_{2})=0.$ (3)
We only consider values of $c$ in $(0,\sqrt{2})$, because if $c$ tends to zero, the amplitude of the waves grows closer to infinity, and for $c=0$, nontrivial solutions do not longer exist [23]. Contrarily, if $c$ approaches $\sqrt{2}$, the solutions become oscillatory as the amplitudes of the waves tend to go to zero [15].
Our goal is to establish the existence of a solution $\tilde{u}$ to (2) on $\Omega$. Numerical simulations of solutions to this equation, obtained via the Mountain Pass Algorithm, are discussed in [15]. A rigorous existence proof for traveling wave solutions to (2) on finite domains, periodic in one direction and satisfying Neumann boundary conditions in the other, was later given in [32]. We combine the techniques and results outlined there with procedures developed in [6] and [7] to obtain an existence proof of a solution that decays to zero as $|x_{1}|\rightarrow\infty$.
Proving the existence of solutions to nonlinear partial differential equations (PDEs) on unbounded domains poses substantial analytical challenges. In particular, the loss of compactness in the resolvent of differential operators on unbounded domains complicates the application of standard existence theorems. In [19], this obstacle is overcome by rigorously controlling the spectrum of the linearization of an approximate solution resulting in methods for proving weak solutions to PDEs of second and fourth order. These methods are demonstrated for the Schrödinger equation on $\mathbb{R}^{2}$. That book also discusses the work of [22], where solutions to Emden’s equation on an unbounded L-shaped domain are proven. Additionally, [33] verifies the existence of a weak solution to the Navier-Stokes equations on a perturbed infinite strip. However, due to limited regularity, these methods do not extend to proving the existence of strong solutions.
In contrast, the previously mentioned papers [6] and [7] provide an approach for proving strong solutions on unbounded domains by introducing a new technique for constructing approximate inverses of PDE operators and proving compactness. In these works, solutions on unbounded domains are proven for equations with polynomial nonlinearities, specifically the Kawahara and Swift-Hohenberg equations. The techniques discussed in these papers need to be adjusted to be able to apply them to our suspension bridge equation that includes an exponential nonlinearity, which will be done in this paper. Our approach is versatile in the sense that it can be extended to other analytic nonlinearities as well.
To deal with nonpolynomial terms such as the exponential term in (2), the method described in [12] can be used to control the aliasing error. In [28], it is demonstrated how to apply these techniques for working with nonpolynomial nonlinearities in one spatial dimension. The generalization to higher dimensions can be found in [32], in which specifically the suspension bridge equation is discussed, but only on bounded domains.
Figure 1: Visualization of an approximate solution $\bar{u}$ to (2) with $c=1.2$. The approximation is truncated to a finite domain in this plot.
Ultimately our goal is to construct an approximate solution $\bar{u}:\Omega\to{\mathbb{R}}$ to (2) and then prove the existence of a true solution to (2) in a vicinity of $\bar{u}$. An illustration of such an approximate solution $\bar{u}$, that we rigorously verify in Theorem 6.1, is depicted in Figure 1. The precise construction of $\bar{u}$ is explained in Section 3.1. The verification of such an approximate solution is done using a computer-assisted proof (CAP), a method that has gained significant interest in the rigorous verification of numerical simulations for dynamical systems over the past few decades. Early implementations of CAPs to the study of elliptic PDEs are seen in [19] and [24], stemming from 1988 and 1992, respectively. Later, advancements in the applications of CAPs to (elliptic) PDEs were made (e.g. [9, 10, 21, 27, 30, 34]). Comprehensive discussions on the role of CAPs in the study of PDEs are outlined in [14] and [20].
In particular, the setting in which we work allows us to utilize a Newton-Kantorovich-type theorem. These type of theorems are widely used in the construction of CAPs for PDEs and various other contexts (see for instance [1, 16, 29]). Specifically, we use a variant of this theorem called the radii polynomial theorem [10]. To compensate for the fact that we are working on an unbounded domain, an additional bound that needs to be computed is introduced, on top of the other bounds mentioned in the radii polynomial theorem. To guarantee the reliability of the computational components of the proof, we employ Julia interval arithmetic using the packages [17] and [26].
The solutions obtained in the CAPs will also be analyzed for stability. This part of the work is based on the procedure outlined in [18], where the authors investigate orbital stability of one-dimensional traveling wave solutions to the suspension bridge equation. Their analysis employs energy methods and spectral analysis to establish the stability results, supported by computer-assisted techniques. In two dimensions, to our knowledge, only numerical results concerning stability are available so far, which can be found in [15]. To rigorously handle the spectral properties of the higher-dimensional equation, we follow the framework proposed in [4]. Using this approach, we are able to conclude that the solution corresponding to the approximation in Figure 1 is orbitally stable.
The paper is organized as follows. We begin by introducing notation and formulating the zero-finding problem. In Section 3, we present the Newton-Kantorovich theorem, which forms the foundation of our existence results. Moreover, we describe the construction of approximate solutions and provide more computational details. Section 4 presents the bounds required to apply the existence theorem. The set-up for the stability analysis is given in Section 5. Finally, the proof of the solution corresponding to the approximation shown in Figure 1 is given in Section 6, along with two additional existence results. All three proofs include verifications of orbital (in)stability.
2 Set-up
In this section, we introduce notation that will be used in the rest of the paper. Part of the notation is borrowed from [6, 7] and adjusted to our setting of the suspension bridge equation.
2.1 Zero-finding problem for functions
Let $\mathbb{L}$ be the linear differential operator defined as
$\mathbb{L}=\Delta^{2}+c^{2}\partial_{x_{1}}^{2}+I$
where $I$ is the identity and $\Delta$ is the Laplace operator. Moreover, let $\mathbb{G}$ be defined as $\mathbb{G}(u)=e^{u}-u-1$. Then, (2) is equivalent to a zero finding problem $\mathbb{F}(u)=0$ where
$\mathbb{F}(u)=\mathbb{L}u+\mathbb{G}(u)=\Delta^{2}u+c^{2}\partial_{x_{1}}^{2}u+e^{u}-1.$ (4)
Note that by defining $\mathbb{G}$ as such we have that $\mathbb{G}(0)=0$ and $D\mathbb{G}(0)=0$. In particular, we look for a solution to (2) such that $u(x)\to 0$ as $|x_{1}|\to\infty$ and $u$ satisfies Neumann boundary conditions (3). Taking advantage of the symmetries $u(x_{1},x_{2})=u(-x_{1},x_{2})=u(x_{1},-x_{2})$, we can write $u$ as a cosine series instead of a general Fourier series. As a result, we choose the following ansatz
$u(x_{1},x_{2})=u_{0}(x_{1})+2\sum_{n_{2}=1}^{\infty}u_{n_{2}}(x_{1})\cos\left(2\pi\tilde{n}_{2}x_{2}\right)~~\text{ with }\tilde{n}_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{n_{2}}{2d_{2}},$ (5)
and where $u_{n_{2}}:\mathbb{R}\to\mathbb{R}$ and $u_{n_{2}}$ is even. Corresponding to (5), we define $L^{2}_{e}$ as the following Hilbert space of $L^{2}$ even functions on the strip $\Omega$:
$L^{2}_{e}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{u(x_{1},x_{2})=u_{0}(x_{1})+2\sum_{n_{2}\in\mathbb{N}}u_{n_{2}}(x_{1})\cos\left(2\pi\tilde{n}_{2}x_{2}\right),~u_{n_{2}}(x)=u_{n_{2}}(-x)\text{ and }\|u\|_{2}<\infty\right\}$
where
$\|u\|_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\|u_{0}\|_{L^{2}({\mathbb{R}})}^{2}+2\sum_{n_{2}\in\mathbb{N}}\|u_{n_{2}}\|_{L^{2}({\mathbb{R}})}^{2}\right)^{\frac{1}{2}}.$
Now, given $u\in L^{2}_{e}$ regular enough, we have that
$\mathbb{L}u(x)=\mathbb{L}_{0}u_{0}(x_{1})+2\sum_{n_{2}\in\mathbb{N}}\mathbb{L}_{n_{2}}u_{n_{2}}(x_{1})\cos\left(2\pi\tilde{n}_{2}x_{2}\right)$
for all $x\in\Omega$, where
$\mathbb{L}_{n_{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\frac{n_{2}^{2}\pi^{2}}{d_{2}^{2}}+\partial_{x_{1}}^{2}\right)^{2}+c^{2}\partial_{x_{1}}^{2}+I$
for all $n_{2}\in\mathbb{N}_{0}.$ In particular, note that for $c\in(0,\sqrt{2})$, $\mathbb{L}_{n_{2}}:H^{4}({\mathbb{R}})\to L^{2}({\mathbb{R}})$ has a bounded inverse (see Lemma 7.1). Consequently, similarly as what was achieved in [6] and [7], we can define the Hilbert space $\mathcal{H}$ as follows:
$\mathcal{H}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{u\in L^{2}_{e},~\|u\|_{\mathcal{H}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\mathbb{L}u\|_{2}<\infty\}.$
In addition, we see that
$\|u\|_{\mathcal{H}}=\left(\|\mathbb{L}_{0}u_{0}\|_{2}^{2}+2\sum_{n_{2}=1}^{\infty}\|\mathbb{L}_{n_{2}}u_{n_{2}}\|_{2}^{2}\right)^{\frac{1}{2}}.$
Using the above, we have that $\mathbb{L}:\mathcal{H}\to L^{2}_{e}$ is an isometric isomorphism. Moreover, we define $l:{\mathbb{R}}^{2}\to{\mathbb{R}}$ as the symbol of $\mathbb{L}$ and $l_{n_{2}}:{\mathbb{R}}\to{\mathbb{R}}$ as the one of $\mathbb{L}_{n_{2}}$ for all $n_{2}\in\mathbb{N}_{0}$. Specifically, we have that
$l(\xi)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|2\pi\xi|^{4}-c^{2}(2\pi\xi_{1})^{2}+1~~\text{ and }~~l_{n_{2}}(\xi_{1})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(2\pi)^{4}(\xi_{1}^{2}+\tilde{n}_{2}^{2})^{2}-c^{2}(2\pi\xi_{1})^{2}+1$ (6)
where we recall that $\tilde{n}_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{n_{2}}{2d_{2}}$.
Now, it remains to prove that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is well defined in order to set-up (2) on $\mathcal{H}$. For that purpose, we use the following result:
Lemma 2.1.
Let $\kappa_{1},\kappa_{2}$ be defined as follows:
$\kappa_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{1-\frac{c^{4}}{4}}~\text{ and }~\kappa_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\left\|\frac{1}{l_{0}}\right\|_{L^{2}({\mathbb{R}})}^{2}+2\sum_{n_{2}\in\mathbb{N}}\left\|\frac{1}{l_{n_{2}}}\right\|_{L^{2}({\mathbb{R}})}^{2}\right)^{\frac{1}{2}}.$ (7)
Then, for all $u,v\in\mathcal{H}$, we have that
$\|(e^{u}-1)v\|_{2}\leq\kappa_{1}(e^{\kappa_{2}\|u\|_{\mathcal{H}}}-1)\|v\|_{\mathcal{H}}.$
Proof.
First, we notice that
$\|(e^{u}-1)v\|_{2}\leq\|e^{u}-1\|_{\infty}\|v\|_{2}.$
Then, we have
$\|e^{u}-1\|_{\infty}\leq e^{\|u\|_{\infty}}-1.$
It remains to prove that $\|u\|_{\infty}\leq\kappa_{2}\|u\|_{\mathcal{H}}$ and $\|v\|_{2}\leq\kappa_{1}\|v\|_{\mathcal{H}}$. Using the proof of Lemma 2.1 in [6], we obtain that
$\|u\|_{\infty}\leq\left(\left\|\frac{1}{l_{0}}\right\|_{L^{2}({\mathbb{R}})}^{2}+2\sum_{n_{2}\in\mathbb{N}}\left\|\frac{1}{l_{n_{2}}}\right\|_{L^{2}({\mathbb{R}})}^{2}\right)^{\frac{1}{2}}\|u\|_{\mathcal{H}}\text{ and }\|v\|_{2}\leq\max_{n_{2}\in\mathbb{N}_{0}}\sup_{\xi_{1}\in{\mathbb{R}}}\left|\frac{1}{l_{n_{2}}(\xi_{1})}\right|\|v\|_{\mathcal{H}}.$
We conclude the proof using (57) in Lemma 7.1. ∎
Remark 2.2.
In practice, for the computation of $\kappa_{2}$, we use Riemann summations to bound the summation up to finitely many $n_{2}$. In fact, this can be achieved using rigorous numerics (cf. LA: Reference code). To estimate the tail of the summation, we use (58) in Lemma 7.1.
Using the above lemma, we have that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is a smooth operator. Accordingly, we investigate the solution of the following zero-finding problem
$\mathbb{F}(u)=0~~\text{ for }u\in\mathcal{H},$ (8)
where $\mathbb{F}:\mathcal{H}\rightarrow L^{2}_{e}$.
2.2 Fourier series representation
As described in [6], existence proofs of localized solutions can be achieved thanks to approximate objects on Fourier coefficients. Following such an approach, we introduce notation corresponding to objects on Fourier coefficients, related to the ones introduced in the previous section. Let $d_{1}>0$ and define $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d_{1},d_{1})\times(-d_{2},d_{2})$. $\Omega_{0}\subset\Omega$ is the bounded domain on which our approximate solution is constructed. We let $\mathbb{N}_{0}^{2}=(\mathbb{N}\cup\{0\})^{2}$ and we define
[…]
for all $n\in\mathbb{N}_{0}^{2}$. The coefficients $\alpha_{n}$ appear when switching between the cosine series representation and the exponential Fourier series one. Let $\ell^{p}$ denote the Lebesgue space for sequences defined as
$\ell^{p}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{U=(U_{n})_{n\in\mathbb{N}_{0}^{2}}:~\|U\|_{p}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left(\sum_{n\in\mathbb{N}_{0}^{2}}\alpha_{n}|U_{n}|^{p}\right)^{\frac{1}{p}}<\infty\right\}.$
Then, operators $\mathbb{L}$ and $\mathbb{G}$ have a Fourier coefficients representation, when being considered on $\Omega_{0}$ with periodic boundary conditions. Indeed, defining $\tilde{n}=(\tilde{n}_{1},\tilde{n}_{2})=\left(\frac{n_{1}}{2d_{1}},\frac{n_{2}}{2d_{2}}\right)$, we introduce $L$, the Fourier coefficients representation of $\mathbb{L}$, as
$(LU)_{n}=l(\tilde{n})U_{n}\text{ for all }n\in\mathbb{N}_{0}^{2}$
and all $U\in\ell^{2}$. Then, we define UVU*V as the discrete convolution for sequences indexed on $\mathbb{N}_{0}^{2}$, given as
$(U\ast V)_{n}=\sum_{m\in\mathbb{Z}^{2}}U_{|m|}V_{|n-m|}$
for $n\in\mathbb{N}_{0}^{2}$. Note that Young’s convolution inequality holds true for this convolution product, i.e., for $U\in\ell^{2}$ and $V\in\ell^{1}$, we have that
$\|U*V\|_{2}\leq\|U\|_{2}\|V\|_{1}.$ (10)
Moreover, given $U\in\ell^{1}$, we define $e^{U}$ as
$e^{U}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\sum_{k=0}^{\infty}\frac{U^{k}}{k!},\vskip-8.5359pt$
where $U^{k}=\overbrace{U*U*\dots*U}^{k\text{ times}}$. In particular, $e^{U}\in\ell^{1}$ for all $U\in\ell^{1}$. Consequently, we define $G$ as the equivalent of $\mathbb{G}$ as
$G(U)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}e^{U}-U-e_{0},$
where $(e_{0})_{n}=1$ if $n=(0,0)$ and $(e_{0})_{n}=0$ for all $n\in\left(\mathbb{N}\setminus\{0\}\right)^{2}$. Finally, we define the zero finding counterpart $F$ on $\Omega_{0}$ as
$F(U)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}LU+G(U).$
Now, as in [6] and [7], we introduce operators which allow us to switch from $L^{2}_{e}$ to $\ell^{2}$ and vice versa. Therefore, we define $\gamma:L^{2}_{e}\rightarrow\ell^{2}$ and $\gamma^{\dagger}:\ell^{2}\rightarrow L^{2}_{e}$ such that
$\left((\gamma(u)\right)_{n}$ $=\frac{1}{|\Omega_{0}|}\int_{\Omega_{0}}u(x_{1},x_{2})e^{-2\pi i\tilde{n}\cdot x}dx\quad\text{ for }n\in\mathbb{N}_{0}^{2},$
$\gamma^{\dagger}(U)(x)$ $=\mathbb{1}_{\Omega_{0}}(x)\sum_{n\in\mathbb{N}_{0}^{2}}\alpha_{n}U_{n}\cos\left(2\pi\tilde{n}_{1}x_{1}\right)\cos\left(2\pi\tilde{n}_{2}x_{2}\right),$
where $\mathbb{1}_{\Omega_{0}}$ denotes the characteristic function on $\Omega_{0}$.
Now, given $X$ a Banach space, we denote $\mathcal{B}(X)$ as the set of bounded linear operators on $X$. Then, we define $\Gamma:\mathcal{B}(L^{2}_{e})\to\mathcal{B}(\ell^{2})$ and $\Gamma^{\dagger}:\mathcal{B}(\ell^{2})\to\mathcal{B}(L^{2}_{e})$ as
$\Gamma(\mathbb{H})=\gamma\mathbb{H}\gamma^{\dagger}~\text{ and }~\Gamma^{\dagger}(H)=\gamma^{\dagger}H\gamma$ (11)
for all bounded linear operators $\mathbb{H}\in\mathcal{B}(L^{2}_{e})$ and $H\in\mathcal{B}(\ell^{2})$. The map $\Gamma^{\dagger}$ will be useful when constructing operators on $\mathcal{B}(L^{2}_{e})$ thanks to operators on Fourier coefficients (that is, operators on $\mathcal{B}(\ell^{2})$). Furthermore, we introduce the norm $\|\cdot\|_{2,\mathcal{H}}$ to denote the operator norm for bounded linear operators on $L^{2}_{e}\rightarrow\mathcal{H}$.
Finally, given $u\in L^{\infty},U\in\ell^{1}$, we introduce notation to denote the linear multiplication operator $\mathbb{M}_{u}$ related to the function $u$ and the linear discrete convolution operator $\mathbb{M}_{U}$ related to the sequence $U$ as
$\mathbb{M}_{u}:L^{2}_{e}\to L^{2}_{e}:{v}\mapsto\mathbb{M}_{u}v=uv~\text{ and }~\mathbb{M}_{U}:\ell^{2}\to\ell^{2}:V\mapsto\mathbb{M}_{U}V=U\ast V.$ (12)
3 Constructive proof of existence
In this section, we present our approach for proving constructively the existence of solutions to (2). Our approach is based on a Newton-Kantorovich theorem. More specifically, we construct a fixed point operator for which fixed points correspond to zeros of $\mathbb{F}$ (8).
First, we construct an approximate solution $\bar{u}\in\mathcal{H}$ thanks to its Fourier coefficients representation. The technical construction of $\bar{u}$ is described in Section 3.1. Given $\bar{u}$, we then construct $\mathbb{A}:L^{2}_{e}\to\mathcal{H}$ as an approximate inverse for the Fréchet derivative $D\mathbb{F}(\bar{u})$. Then, our goal is to determine a radius $r>0$ such that $\mathbb{T}$, given as
$\mathbb{T}(u)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}u-\mathbb{A}\mathbb{F}(u),$
is well defined on $\overline{B_{r}(\bar{u})}\to\overline{B_{r}(\bar{u})}$ and contracting (where $B_{r}(\bar{u})$ is the open ball of $\mathcal{H}$ centered at $\bar{u}$ and of radius $r$). The Banach fixed point theorem then provides the existence of a unique zero $\tilde{u}\in\mathcal{H}$ in $\overline{B_{r}(\bar{u})}.$
In order to verify the existence of such a radius $r$, we make use of the following result. Its proof is for instance given in [6].
Theorem 3.1.
Let $\mathbb{F}:\mathcal{H}\to L^{2}_{e}$ be given as in (4) and let $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}$ be a bounded linear operator. Assume that $\mathcal{Y}_{0},\mathcal{Z}_{1}\in(0,\infty)$ and let $\mathcal{Z}_{2}:(0,\infty)\to[0,\infty)$ satisfy
$\|\mathbb{A}\mathbb{F}(\bar{u})\|_{\mathcal{H}}$ $\leq\mathcal{Y}_{0},$ (13)
$\|I-\mathbb{A}D\mathbb{F}(\bar{u})\|_{\mathcal{H}}$ $\leq\mathcal{Z}_{1},$ (14)
$\|\mathbb{A}(D\mathbb{F}(\bar{u}+h)-D\mathbb{F}(\bar{u}))\|_{\mathcal{H}}$ $\leq\mathcal{Z}_{2}(r)r\quad\text{ for all }h\in\overline{B_{r}(0)}\subset\mathcal{H}\text{ and all }r>0.$ (15)
If there exists $r>0$ such that
$\frac{1}{2}\mathcal{Z}_{2}(r)r^{2}-(1-\mathcal{Z}_{1})r+\mathcal{Y}_{0}<0\text{ and }\mathcal{Z}_{1}+\mathcal{Z}_{2}(r)r<1,$ (16)
then there exists a unique $\tilde{u}\in\overline{B_{r}(\bar{u})}\subset\mathcal{H}$ for which $\mathbb{F}(\tilde{u})=0$.
Observe that we have transformed our problem into computing explicit values for the bounds $\mathcal{Y}_{0},\mathcal{Z}_{1},\mathcal{Z}_{2}$ of the previous theorem. Then, a value for the radius $r>0$ is obtained thanks to (16).
The core of our approach becomes determining $\bar{u}\in\mathcal{H}$ and $\mathbb{A}:L^{2}_{e}\to\mathcal{H}$ such that the bounds of Theorem 3.1 can be computed explicitly. We describe the construction of these objects in the next sections.
3.1 Construction of an approximate solution
First, we demonstrate how to construct the approximate solution $\bar{u}\in\mathcal{H}$. Our approach is based on that presented in [5, 6, 7]. We construct $\bar{u}$ using a cosine-cosine series representation on ${\Omega_{0}}=(-d_{1},d_{1})\times(-d_{2},d_{2})$ as
$\bar{u}(x)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{1}_{\Omega_{0}}(x)\sum_{n\in\mathbb{N}_{0}^{2}}\alpha_{n}\bar{U}_{n}\cos\left(2\pi\tilde{n}_{1}x_{1}\right)\cos\left(2\pi\tilde{n}_{2}x_{2}\right)$ (17)
[…]
where $I_{\mathcal{N}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{n\in\mathbb{N}_{0}^{2}:0\leq n_{1}\leq\mathcal{N}_{1},0\leq n_{2}\leq\mathcal{N}_{2}\}$.
Let us fix $N^{0}=(N^{0}_{1},N^{0}_{2})\in\mathbb{N}^{2}$ to be the numerical truncation size of $\bar{U}$. Then, using (18), we assume that $\bar{U}=\pi^{N^{0}}\bar{U}$. That is, $\bar{U}$ has a vector representation of size $(N_{1}^{0}+1)(N_{2}^{0}+1).$
Now, note that by construction $\bar{u}\in L^{2}_{e}$ but we do not necessarily have $\bar{u}\in\mathcal{H}.$ Indeed, $\bar{u}$ might not be smooth at $x_{1}=\pm d_{1}$. By periodicity of $\bar{u}$ on $\Omega_{0}$}, this is equivalent for asking smoothness only at $d_{1}$. Now, in order to obtain the required regularity for $\mathcal{H}$, we need
$\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0\text{ for all }x_{2}\in(-d_{2},d_{2})\text{ and all }k\in\{0,1,2,3\}.$
Note that since $\bar{u}$ has a cosine-cosine representation, we readily have $\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0$ for all $x_{2}\in(-d_{2},d_{2})$ and all $k\in\{1,3\}$. Consequently, it remains to ensure that
$\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0\text{ for all }x_{2}\in(-d_{2},d_{2})\text{ and all }k\in\{0,2\}.$ (19)
Note that since $\bar{u}$ is represented thanks to $\bar{U}$ on $\Omega_{0}$}, (19) can be translated as equations on $\bar{U}$ :
$\sum_{n_{1}=-N^{0}_{1}}^{N^{0}_{1}}(-1)^{n_{1}}(2\pi\tilde{n}_{1})^{k}\bar{U}_{n_{1},n_{2}}=0\text{ for all }n_{2}\in-N^{0}_{2},\dots,N^{0}_{2}\text{ and all }k\in\{0,2\},$
where we have used that $\bar{U}=\pi^{N}\bar{U}$. Note that this is equivalent to having $\bar{U}\in Ker(\mathcal{T})$, where
$\mathcal{T}\bar{U}=\begin{pmatrix}\mathcal{T}_{0}\bar{U}\\ \mathcal{T}_{2}\bar{U}\end{pmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{pmatrix}\left(\sum_{n_{1}=-N^{0}_{1}}^{N^{0}_{1}}(-1)^{n_{1}}\bar{U}_{n_{1},n_{2}}\right)_{n_{2}\in-N^{0}_{2},\dots,N^{0}_{2}}\\ \left(\sum_{n_{1}=-N^{0}_{1}}^{N^{0}_{1}}(-1)^{n_{1}}(2\pi\tilde{n}_{1})^{2}\bar{U}_{n_{1},n_{2}}\right)_{n_{2}\in-N^{0}_{2},\dots,N^{0}_{2}}\end{pmatrix}$.
Consequently, $\bar{U}$ needs to be projected into the kernel of the matrix $\mathcal{T}$. Such an operation has been described in [6, 7] and is implemented in [31].
Overall, we ensure that $\bar{U}\in Ker(\mathcal{T})$ and obtain that
$\bar{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}\left(\bar{U}\right)\in\mathcal{H}~\text{ and }~\bar{U}=\pi^{N}\bar{U}.$ (20)
[…]
Now that $\bar{u}$ has been constructed, one of the major difficulties in the treatment of (2) is the evaluation of the nonpolynomial nonlinearity $e^{\bar{u}}$. Indeed, in order to evaluate the bound $\mathcal{Y}_{0}$ (13), we need to be able to compute rigorously such a quantity. This is the objective of the next section.
3.2 Computing the nonlinearity
The existence proof requires the computation of the nonlinear part ${G}(U)=e^{U}-U-e_{0}$. However, the Fourier coefficients corresponding to the exponential function cannot be determined exactly, as there are infinitely many nonzero Fourier coefficients. Moreover, even the computation of a finite number of Fourier coefficients cannot be performed exactly due to the nonpolynomial nature of the nonlinearity. For small indices, we approximate these coefficients using discrete Fourier transforms and establish an upper bound on the associated error, referred to as the aliasing error. For larger indices, we use a general bound, which is only suitable for the tail part because of the decay of the bound. To compute the Fourier coefficients of ${G}(\bar{U})$, we first determine the Fourier coefficients $e^{\bar{U}}$ corresponding to the function $e^{\bar{u}}$. Following the procedure outlined in Section 3 of [32], we obtain the expression:
$e^{\bar{U}}\in\begin{cases}{(e^{\bar{U}})}^{\mathrm{FFT}}_{n}+{C}[-\varepsilon_{n},\varepsilon_{n}]&\text{ if }0\leq n_{1}\leq N^{0}_{1},0\leq n_{2}\leq N^{0}_{2},\\ [-\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}},\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}}]&\text{ otherwise}\end{cases}$ (21)
where ${(e^{\bar{U}})}^{\mathrm{FFT}}$ is the numerical approximation of $e^{\bar{U}}$ calculated using discrete Fourier transforms, $C>0$, $\nu_{1},\nu_{2}>1$ and $\varepsilon_{n}>0$ for all $n$. The procedure of computing $C$ can be found in [32]. The parameters $\nu_{1},\nu_{2}$ are related to the domain of analyticity, i.e., $\bar{U}$ is analytic on a strip $\{z\in\mathbb{C}^{2}:\mathrm{Im}(z_{1})<\rho_{1},\mathrm{Im}(z_{2})<\rho_{2}\}$ and $\nu_{1}=e^{\rho_{1}-\overline{\varepsilon}_{1}}$, $\nu_{2}=e^{\rho_{2}-\overline{\varepsilon}_{2}}$ for small $\overline{\varepsilon}_{1},\overline{\varepsilon}_{2}>0$ (cf. Section 3.1 in [32]). Note that ${G}(\bar{U})$ is analytic for $\bar{U}$ analytic.
Moreover, let $N^{\mathrm{FFT}}$ denote how many Fourier modes are considered when doing Fourier transformations such that $N^{\mathrm{FFT}}_{1}$ and $N^{\mathrm{FFT}}_{2}$ are powers of $2$ and $N^{\mathrm{FFT}}\gg N^{0}$. Then we have
$\varepsilon_{n}:=\frac{2\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}\left(\nu_{1}^{-2N_{1}^{\mathrm{FFT}}}+\nu_{2}^{-2N_{2}^{\mathrm{FFT}}}\right)}{\left(1-\nu_{1}^{-2N_{1}^{\mathrm{FFT}}}\right)\left(1-\nu_{2}^{-2N_{2}^{\mathrm{FFT}}}\right)}.$
Remark 3.2.
Note that the support of $\bar{u}$ is on $\overline{\Omega_{0}}$, which is a finite domain. As a consequence, $\mathbb{G}(\bar{u})$ also has its support on $\overline{\Omega_{0}}$. Therefore, the actual representation of the nonlinear term is on the infinite strip, even though Fourier transformations are carried out on finite domains only.
3.3 Approximate inverse
In order to be able to apply Theorem 3.1, it remains to construct an approximate inverse $\mathbb{A}$ for $D\mathbb{F}(\bar{u})$. For this purpose, we recall the construction established in [6]. Let $N=(N_{1},N_{2})\in\mathbb{N}^{2}$ be the numerical truncation size of our operators. In practice, we choose $N_{1}\leq N^{0}_{1}$ and $N_{2}\leq N^{0}_{2}$, where $N^{0}$ is the size of our sequences.
Numerically, we construct a “matrix” $B^{N}$ approximating the inverse of $\pi^{N}DF(\bar{U})L^{-1}\pi^{N}$. In particular, $B^{N}$ is finite-dimensional in the sense that $B^{N}=\pi^{N}B^{N}\pi^{N}$. Then, using (11), we define $\mathbb{B}:L^{2}_{e}\to L^{2}_{e}$ as
$\mathbb{B}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{1}_{\Omega\backslash\Omega_{0}}+\Gamma^{\dagger}(\pi_{N}+B^{N}).$
Here, $\mathbb{1}_{\Omega\backslash\Omega_{0}}$ has to be understood as a multiplication operator by the characteristic function $\mathbb{1}_{\Omega\backslash\Omega_{0}}$. In particular, $\mathbb{B}$ approximates the inverse of $D\mathbb{F}(\bar{u})\mathbb{L}^{-1}.$ Finally, we define $\mathbb{A}$ as
$\mathbb{A}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{L}^{-1}\mathbb{B}=\mathbb{L}^{-1}\left(\mathbb{1}_{\Omega\backslash\Omega_{0}}+\Gamma^{\dagger}(\pi_{N}+B^{N})\right):L^{2}_{e}\to\mathcal{H}.$ (22)
Note that $\mathbb{A}$ is well defined as $\mathbb{L}:\mathcal{H}\to L^{2}_{e}$ is an isometric isomorphism. Computing the bound $\mathcal{Z}_{1}$ defined in (14), we will show that (22) is an accurate choice for an approximate inverse. In particular, it is well suited for our analysis since it allows us to compute the bounds of Theorem 3.1 explicitly. This will be the goal of the next section. We can already recall a result from [6] providing the operator norm of $\mathbb{A}$ thanks to the norm of the matrix $B^{N}$:
$\|\mathbb{A}\|_{2,\mathcal{H}}=\|\mathbb{B}\|_{2}=\max\{1,\|B^{N}\|_{2}\}.$ (23)
4 Computation of the bounds
In this section, we justify our choices for the approximate objects constructed above by providing explicit formulas for the bounds of Theorem 3.1.
Lemma 4.1 (Bound $\mathcal{Y}_{0}$).
Let $\mathcal{Y}_{0}$ be satisfying
$\mathcal{Y}_{0}\geq\sqrt{2d_{1}}$ $\Bigg(\|B^{N}F(\bar{U})\|_{2}^{2}+\|(\pi^{N^{0}}-\pi^{N})(L\bar{U}+G(\bar{U}))\|_{2}^{2}$
$+C^{2}\frac{2\nu_{1}^{-2N^{0}_{1}-2}(1+\nu_{2}^{-2})+2\nu_{2}^{-2N^{0}_{2}-2}(1+\nu_{1}^{-2})-4\nu_{1}^{-2N_{1}^{0}-2}\nu_{2}^{-2N_{2}^{0}-2}}{(1-\nu_{1}^{-2})(1-\nu_{2}^{-2})}\Bigg)^{\frac{1}{2}}.$
Then $\mathcal{Y}_{0}$ satisfies (13).
Proof.
As $\mathrm{supp}(\bar{u})\subset\overline{\Omega_{0}}$, we have that $\mathrm{supp}(\mathbb{G}(\bar{u}))\subset\overline{\Omega_{0}}$ since $e^{u}-u-1=\sum_{k=2}^{\infty}\frac{u^{k}}{k!}$. Hence, we can apply Parseval’s identity and follow the proof of Lemma 4.11 in [6] to obtain
$\|\mathbb{A}\mathbb{F}(\bar{u})\|_{\mathcal{H}}=\|\mathbb{L}\mathbb{A}\mathbb{F}(\bar{u})\|_{2}=\|\mathbb{B}\mathbb{F}(\bar{u})\|_{2}=\sqrt{2d_{1}}\|BF(\bar{U})\|_{2}$
where
$\|BF(\bar{U})\|_{2}^{2}=\|B^{N}F(\bar{U})\|_{2}^{2}+\|(\pi^{N^{0}}-\pi^{N})L\bar{U}+\pi_{N}G(\bar{U})\|_{2}^{2}.$ (24)
Recall that $N^{0}$ corresponds to the truncation dimension for sequences and $N$ for operators. Note that the nonlinear part of $F(\bar{U})$ in $\|B^{N}F(\bar{U})\|_{2}^{2}$ can be calculated using (21).
For the second term in the right-hand side of (24) we write
$\|(\pi^{N^{0}}-\pi^{N})L\bar{U}+\pi_{N}G(\bar{U})\|_{2}^{2}$ $\leq\|(\pi^{N^{0}}-\pi^{N})(L\bar{U}+G(\bar{U}))\|_{2}^{2}$
$+\sum_{n\notin I_{N^{0}}}\alpha_{n}\frac{C^{2}}{\nu_{1}^{2|n_{1}|}\nu_{2}^{2|n_{2}|}}.$
The last summation can be calculated using a geometric series argument as was done in Lemma 3.4 of [32]. Hence,
$\sum_{n\notin I_{N^{0}}}$ $\alpha_{n}\frac{C^{2}}{\nu_{1}^{2|n_{1}|}\nu_{2}^{2|n_{2}|}}$
$=C^{2}\frac{2\nu_{1}^{-2N^{0}_{1}-2}(1+\nu_{2}^{-2})+2\nu_{2}^{-2N^{0}_{2}-2}(1+\nu_{1}^{-2})-4\nu_{1}^{-2N_{1}^{0}-2}\nu_{2}^{-2N_{2}^{0}-2}}{(1-\nu_{1}^{-2})(1-\nu_{2}^{-2})}.$
Remark 4.2.
Note that for the second term in the $\mathcal{Y}_{0}$-bound, we can use (21) to state the interval that contains this term as follows:
$\|(\pi^{N^{0}}-\pi^{N})(L\bar{U}+G(\bar{U}))\|_{2}^{2}\in\sum_{\begin{subarray}{c}n\in I_{N^{0}}\\ n\notin I_{N}\end{subarray}}\alpha_{n}\left(l(\tilde{n})(\bar{U})_{n}+(e^{\bar{U}})^{\mathrm{FFT}}+C[-\varepsilon_{n},\varepsilon_{n}]\right)^{2}.$
The right-hand side is what is used in the (interval arithmetic) computation of the $\mathcal{Y}_{0}$-bound in [31].
We now turn to the computation of $\mathcal{Z}_{2}(r)$.
Lemma 4.3 (Bound $\mathcal{Z}_{2}$).
Let $r>0$ and let $\mathcal{Z}_{2}(r)$ be satisfying
$\mathcal{Z}_{2}(r)\geq\kappa_{1}\frac{e^{\kappa_{2}r}-1}{r}\max\left\{\|{B}e^{\mathbb{M}_{\bar{U}}}\|_{2},1\right\},$
where $\kappa_{1}$ and $\kappa_{2}$ are given in Lemma 2.1. Then $\mathcal{Z}_{2}(r)$ satisfies (15).
Proof.
Let $h\in\overline{B_{r}(0)}$. Since $\|u\|_{\mathcal{H}}=\|\mathbb{L}u\|_{{2}}$ and $\mathbb{A}=\mathbb{L}^{-1}\mathbb{B}$ by definition, we have that
$\|\mathbb{A}(D\mathbb{F}(\bar{u}+h)-D\mathbb{F}(\bar{u}))\|_{\mathcal{H}}$ $=\|\mathbb{L}\mathbb{A}(D\mathbb{F}(\bar{u}+h)-D\mathbb{F}(\bar{u}))\|_{\mathcal{H},{2}}$
$=\|\mathbb{B}(e^{\mathbb{M}_{\bar{u}+h}}-e^{\mathbb{M}_{\bar{u}}})\|_{\mathcal{H},{2}}.$
Then, we have
$\|\mathbb{B}(e^{\mathbb{M}_{\bar{u}+h}}-e^{\mathbb{M}_{\bar{u}}})\|_{\mathcal{H},{2}}$ $\leq\underset{\|w\|_{\mathcal{H}}=1}{\sup}\|\mathbb{B}e^{\mathbb{M}_{\bar{u}}}\|_{2}\|(e^{h}-1)w\|_{L^{2}_{e}}$
where, using (23), we also get
$\|\mathbb{B}e^{\mathbb{M}_{\bar{u}}}\|_{2}\leq\max\left\{\|{B}e^{\mathbb{M}_{\bar{U}}}\|_{2},1\right\}.$
We conclude the proof using Lemma 2.1, from which we deduce that
$\|(e^{h}-1)w\|_{L^{2}_{e}}\leq\kappa_{1}(e^{\kappa_{2}r}-1).$
For the computation of the bound $\mathcal{Z}_{1}$, we first introduce some extra notation for simplicity. Let $\overline{v}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}e^{\bar{u}}-1$ and $\bar{V}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma(\bar{v})$. Doing so, notice that $D\mathbb{G}(\bar{u})$ and $DG(\bar{U})$ can be expressed as
$D\mathbb{G}(\bar{u})=\mathbb{M}_{\bar{v}}\text{ and }DG(\bar{U})=M_{\bar{V}},$
where we have used the notation in (12). Finally, note that $\bar{V}$ is an infinite number of Fourier coefficients, as explained in Section 3.2. In order to exhibit finite-dimensional computations, we define $\bar{V}^{N}$ and $\bar{v}^{N}$ as
$\bar{V}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\pi^{N}\bar{V}\text{ and }\bar{v}^{N}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(\bar{V}^{N}).$
That is, $\bar{V}^{N}$ correspond to the $N$ first modes of $\bar{V}$. In fact, the difference $\bar{V}-\bar{V}^{N}$ can be controlled explicitly as a result of the analysis of Section 3.2. Then, we recall Lemma 3.6 from [7], providing the decomposition of $\mathcal{Z}_{1}$ as a Fourier coefficients computation $Z_{1}$, and a remainder $\mathcal{Z}_{u}$ that corresponds to the unboundedness of the domain $\Omega.$
Lemma 4.4 (Bound $\mathcal{Z}_{1}$).
Let $\mathcal{Z}_{u}$ and $Z_{1}$ be bounds satisfying
$\|\left(\mathbb{L}^{-1}-\Gamma^{\dagger}(L^{-1})\right)\mathbb{M}_{\overline{v}^{N}}\mathbb{B}^{*}\|_{2}$ $\leq\mathcal{Z}_{u},$ (25)
$\|I-(B^{N}+\pi_{N})(I+M_{\bar{V}^{N}}L^{-1})\|_{2}$ $\leq Z_{1}.$
Then defining $\mathcal{Z}_{1}$ as
$\mathcal{Z}_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}Z_{1}+\left(\mathcal{Z}_{u}+\frac{\max\{1,\|B^{N}\|_{2}\}}{1-\frac{c^{4}}{4}}\|\bar{V}-\bar{V}^{N}\|_{1}\right),$ (26)
we get that (14) holds, that is $\|I-\mathbb{A}D\mathbb{F}(\bar{u})\|_{\mathcal{H}}\leq\mathcal{Z}_{1}.$
Proof.
The proof is given in [7]. Note that the difference in the formulas lies in the fact that $\sup_{\xi\in{\mathbb{R}}^{2}}\frac{1}{l(\xi)}=\frac{1}{1-\frac{c^{4}}{4}}$ in the case of (2). ∎
First, we recall the computation of the bound $Z_{1}$, which is given in Lemma 3.9 from [7].
Lemma 4.5.
Let $Z^{N}_{1}$ and $Z_{1}$ be such that
$\left(\|\pi^{N}-B^{N}(I_{d}+M_{\bar{V}^{N}}L^{-1})\pi^{2N}\|_{2}^{2}+\|(\pi^{2N}-\pi^{N})M_{\bar{V}^{N}}L^{-1}\pi^{N}\|_{2}^{2}\right)^{\frac{1}{2}}$ $\leq Z_{1}^{N},$ (27)
$\left((Z_{1}^{N})^{2}+\|\bar{V}^{N}\|_{1}^{2}\max_{n\in\mathbb{N}_{0}^{2}\setminus I^{N}}\frac{1}{|l(\tilde{n})|^{2}}\right)^{\frac{1}{2}}$ $\leq Z_{1}.$
Then we have $\left\|I_{d}-(B^{N}+\pi_{N})\left(I_{d}+M_{\bar{V}^{N}}L^{-1}\right)\right\|_{2}\leq Z_{1}$.
The above quantities are related to finite-dimensional computations, and their computation (with interval arithmetic) is presented in [31]. It remains to compute the bound $\mathcal{Z}_{u}$ satisfying (25).
Following the analysis of [6], we can write $\mathcal{Z}_{u}$ as
$\mathcal{Z}_{u}=\max\left\{1,\|B^{N}\|_{2}\right\}\sqrt{\mathcal{Z}_{u,1}^{2}+\mathcal{Z}_{u,2}^{2}}$
where $\mathcal{Z}_{u,1}$ and $\mathcal{Z}_{u,2}$ are defined as follows:
$\mathcal{Z}_{u,1}$ $=\sup_{\|u\|_{2}=1}\left(\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{0}^{-1}(\bar{v}^{N}u)_{0}\|_{2}^{2}+2\sum_{n_{2}\in\mathbb{N}}\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{n_{2}}^{-1}(\bar{v}^{N}u)_{n_{2}}\|_{2}^{2}\right)^{\frac{1}{2}},$
$\mathcal{Z}_{u,2}$ $=\sup_{\|u\|_{2}=1}\left(\|\mathbb{1}_{(-d_{1},d_{1})}(\mathbb{L}_{0}^{-1}-\Gamma^{\dagger}(L_{0}^{-1}))(\bar{v}^{N}u)_{0}\|_{2}^{2}+2\sum_{n_{2}\in\mathbb{N}}\|\mathbb{1}_{(-d_{1},d_{1})}(\mathbb{L}_{n_{2}}^{-1}-\Gamma^{\dagger}(L_{n_{2}}^{-1}))(\bar{v}^{N}u)_{n_{2}}\|_{2}^{2}\right)^{\frac{1}{2}}.$
In particular, we have that $\mathbb{L}_{n_{2}}^{-1}$ is a convolution operator associated to the function $f_{n_{2}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{F}^{-1}\left(\frac{1}{l_{n_{2}}}\right)$. Indeed, we have that
$\mathbb{L}_{n_{2}}u=\mathcal{F}^{-1}\left(\mathcal{F}(\mathbb{L}_{n_{2}}u)\right)=\mathcal{F}^{-1}\left(\frac{1}{l_{n_{2}}}\hat{u}\right)=f_{n_{2}}*u.$
The results derived in [6] require the computation of constants $C_{n_{2}}>0$ and $a_{n_{2}}>0$ such that
$|f_{n_{2}}(x)|\leq C_{n_{2}}e^{-a_{n_{2}}|x|},~~\text{ for all }x\in{\mathbb{R}}.$ (28)
If we can obtain explicit constants satisfying the above, then [6] provides explicit formulas for the computation of $\mathcal{Z}_{u,1}$ and $\mathcal{Z}_{u,2}$. We derive such a result in the following lemma.
Lemma 4.6.
Let $n_{2}\in\mathbb{N}_{0}$, and let us define $a_{n_{2}}$, $b_{n_{2}}$ and ${C}_{n_{2}}$ as
$a_{n_{2}}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\sqrt{4(1+\tilde{n}_{2}^{2}c^{2})-c^{4}}}{2\sqrt{c^{2}-2\tilde{n}_{2}^{2}+\sqrt{(c^{2}-2\tilde{n}_{2}^{2})^{2}+4(1+c^{2}\tilde{n}_{2}^{2})-c^{4}}}},$ (29)
$b_{n_{2}}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2}\sqrt{c^{2}-2\tilde{n}_{2}^{2}+\sqrt{(c^{2}-2\tilde{n}_{2}^{2})^{2}+4(1+c^{2}\tilde{n}_{2}^{2})-c^{4}}},$ (30)
$C_{n_{2}}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2(a_{n_{2}}^{2}+b_{n_{2}}^{2})^{\frac{1}{2}}\pi\sqrt{4(1+\tilde{n}_{2}^{2}c^{2})-c^{4}}}.$ (31)
Then, we obtain that ${C}_{n_{2}}$ and $a_{n_{2}}$ satisfy (28).
Proof.
Let $n_{2}\in\mathbb{N}_{0}$ and let us compute the inverse Fourier transform of $\frac{1}{l_{n_{2}}}$. We have
$f_{n_{2}}(x)=\int_{{\mathbb{R}}}\frac{1}{l_{n_{2}}(\xi)}e^{2\pi i\xi x}d\xi=\frac{1}{2\pi}\int_{{\mathbb{R}}}\frac{1}{l_{n_{2}}(\frac{\xi}{2\pi})}e^{i\xi x}d\xi.$
Now, notice that $\xi\mapsto l_{n_{2}}(\frac{\xi}{2\pi})$ has four roots given by $\pm\xi_{1}$ and $\pm\overline{\xi_{1}}$ where
$\xi_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}b_{n_{2}}-ia_{n_{2}}.$
In particular, we have that
$\frac{1}{l_{n_{2}}(\frac{\xi}{2\pi})}=\frac{1}{i\sqrt{4(1+\tilde{n}_{2}^{2}c^{2})-c^{4}}}\left(\frac{1}{\xi^{2}-\xi_{1}^{2}}-\frac{1}{\xi^{2}-\overline{\xi_{1}}^{2}}\right).$
Using Section 2.3.14 in [25], we obtain that
$f_{n_{2}}(x)$ $=\frac{1}{2\pi}\frac{1}{i\sqrt{4(1+\tilde{n}_{2}^{2}c^{2})-c^{4}}}\int_{{\mathbb{R}}}\left(\frac{1}{\xi^{2}-\xi_{1}^{2}}-\frac{1}{\xi^{2}-\overline{\xi_{1}}^{2}}\right)e^{i\xi x}d\xi$
$=\frac{1}{4\pi}\frac{1}{i\sqrt{4(1+\tilde{n}_{2}^{2}c^{2})-c^{4}}}\left(\frac{e^{-i\xi_{1}|x|}}{i\xi_{1}}-\frac{e^{-\overline{i\xi_{1}}|x|}}{\overline{i\xi_{1}}}\right)$
for all $x\in{\mathbb{R}}$. Then, we have that
$f_{n_{2}}(x)$ $=\frac{e^{-a_{n_{2}}|x|}}{4(a_{n_{2}}^{2}+b_{n_{2}}^{2})i\pi\sqrt{4(1+\tilde{n}_{2}^{2}c^{2})-c^{4}}}\left(e^{-ib_{n_{2}}|x|}(a_{n_{2}}-ib_{n_{2}})-e^{ib_{n_{2}}|x|}(a_{n_{2}}+ib_{n_{2}})\right)$
$=-\frac{e^{-a_{n_{2}}|x|}}{2(a_{n_{2}}^{2}+b_{n_{2}}^{2})\pi\sqrt{4(1+\tilde{n}_{2}^{2}c^{2})-c^{4}}}\left(a_{n_{2}}\text{sign}(x)\sin(b_{n_{2}}x)+b_{n_{2}}\cos(b_{n_{2}}x)\right).$
Notice that $x\mapsto f_{n_{2}}(x)$ is an even function and that $f_{n_{2}}$ is smooth on $0,\infty)$. In particular, we have that $C_{n_{2}}$ and $a_{n_{2}}$ satisfy (28). ∎
Using the above, we obtain an explicit formula for $\mathcal{Z}_{u}$.
Lemma 4.7.
Let $E\in\ell^{2}$ the sequence of Fourier coefficients given as
$E_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{(C_{n_{2}})^{2}a_{n_{2}}(-1)^{n_{1}}(1-e^{-4a_{n_{2}}d_{1}})}{d_{1}(4a_{n_{2}}^{2}+(2\pi\tilde{n}_{1})^{2})}$ (32)
for all $n\in\mathbb{N}_{0}^{2}$. Moreover, let $a$ and $C(d_{1})$ be defined as
$a\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\inf_{n_{2}\in\mathbb{N}_{0}}a_{n_{2}}~\text{ and }~C(d_{1})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4d_{1}+\frac{4e^{-ad_{1}}}{a(1-e^{-\frac{3}{2}ad_{1}})}+\frac{2}{a(1-e^{-2ad_{1}})}.$
Then, we obtain that
$\mathcal{Z}_{u,1}$ $\leq 2d_{1}(\bar{V}^{N},\bar{V}^{N}*E)_{2},$
$\mathcal{Z}_{u,2}$ $\leq 2d_{1}(\bar{V}^{N},V^{N}*E)_{2}+C(d_{1})e^{-2ad_{1}}(\bar{V}^{N},\bar{V}^{N}*E)_{2},$
and that $\mathcal{Z}_{u}=\max\left\{1,\|B^{N}\|_{2}\right\}\sqrt{\mathcal{Z}_{u,1}^{2}+\mathcal{Z}_{u,2}^{2}}$ satisfies (25).
Proof.
Let $u\in L^{2}_{e}$ such that $\|u\|_{L^{2}_{e}}=1$ and let $v\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\bar{v}^{N}u$. Using the proof of Theorem 3.9 in [6], we have
$\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{n_{2}}^{-1}v_{n_{2}}\|_{2}^{2}\leq C_{n_{2}}^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\left(\int_{-d_{1}}^{d_{1}}e^{-a_{n_{2}}|x-y|}|v_{n_{2}}(x)|dx\right)^{2}dy$ (33)
for all $n_{2}\in\mathbb{N}_{0}$, where we have used (28). Using that $v=\bar{v}^{N}u$, we have
$v_{n_{2}}=\sum_{k\in\mathbb{Z}}\bar{v}^{N}_{|n_{2}-k|}u_{|k|}.$
for all $n_{2}\in\mathbb{N}_{0}$. Going back to (33), we get
$\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{n_{2}}^{-1}v_{n_{2}}\|_{2}^{2}\leq C_{n_{2}}^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\left(\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}e^{-a_{n_{2}}|x-y|}|\bar{v}^{N}_{|n_{2}-k|}(x)u_{|k|}(x)|dx\right)^{2}dy.$
Now, using the Cauchy-Schwarz inequality, we get
$\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{n_{2}}^{-1}v_{n_{2}}\|_{2}^{2}$
$\leq$~ $C_{n_{2}}^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\left(\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}e^{-2a_{n_{2}}|x-y|}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}dx\right)\left(\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}|u_{|k|}(x)|^{2}dx\right)dy$
$\leq$~ $C_{n_{2}}^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}e^{-2a_{n_{2}}|x-y|}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}dxdy$ (34)
since $\|u\|_{2}=1.$ Now, using Fubini’s theorem, we have
$\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}e^{-2a_{n_{2}}|x-y|}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}dxdy=\int_{-d_{1}}^{d_{1}}\sum_{k\in\mathbb{Z}}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\ e^{-2a_{n_{2}}|x-y|}dydx.$
Using the proof of Theorem 3.9 from [6] again, we get
$\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\ e^{-2a_{n_{2}}|x-y|}dy=\frac{e^{-2a_{n_{2}}d_{1}}\cosh(2a_{n_{2}}x)}{a_{n_{2}}}.$
In particular, using Lemma 4.6 in [3], we obtain that $x\to\frac{e^{-2a_{n_{2}}d_{1}}\cosh(2a_{n_{2}}x)}{a_{n_{2}}}$ has a Fourier coefficients representation on $(-d_{1},d_{1})$ given by the sequence $\tilde{E}$, where
$\tilde{E}_{n_{1}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{a_{n_{2}}(-1)^{n_{1}}(1-e^{-4a_{n_{2}}d_{1}})}{d_{1}(4a_{n_{2}}^{2}+(2\pi\tilde{n}_{1})^{2})}$
for all $n_{1}\in\mathbb{N}_{0}$ and all $n_{2}\in\mathbb{N}_{0}$. Using Parseval’s identity, this implies that
$\mathcal{Z}_{u,1}^{2}$ $\leq\sum_{k\in\mathbb{Z}}\sum_{n_{2}\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}\frac{C_{|n_{2}|}^{2}e^{-2a_{n_{2}}d_{1}}\cosh(2a_{n_{2}}x)}{a_{n_{2}}}dx$
$=2d_{1}(\bar{V}^{N},\bar{V}^{N}*E)_{2}.$ (35)
This concludes the proof for $\mathcal{Z}_{u,1}$.
For the computation of $\mathcal{Z}_{u,2}$ we use the formulas provided in [6]. Indeed, using the proof of Theorem 3.9, we can define $g_{n}$ as
$g_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2d_{1}}\int_{-d_{1}}^{d_{1}}\mathbb{L}_{n_{2}}v_{n_{2}}(x)e^{-2\pi i\tilde{n}_{1}x}dx$
for all $n\in\mathbb{Z}^{2}$, where $v_{n_{2}}=v_{-n_{2}}$ if $n_{2}<0$. Using the proof of Theorem 3.9 in [6], we have that
$\mathcal{Z}_{u,2}^{2}=2d_{1}\sum_{n\in\mathbb{Z}^{2}}|g_{n}|^{2}=\int_{{\mathbb{R}}\setminus((-d_{1},d_{1})\cup(-d_{1},d_{1})+2d_{1}n_{1})}\mathbb{L}_{n_{2}}v_{n_{2}}(x)\mathbb{L}_{n_{2}}v_{n_{2}}(x-2d_{1}n_{1})dx.$
In particular, note that for $n_{1}=0$, we have
$\sum_{n_{2}\in\mathbb{Z}}\int_{{\mathbb{R}}\setminus((-d_{1},d_{1})\cup(-d_{1},d_{1})+2d_{1}n_{1})}\mathbb{L}_{n_{2}}v_{n_{2}}(x)\mathbb{L}_{n_{2}}v_{n_{2}}(x-2d_{1}n_{1})dx=\sum_{n_{2}\in\mathbb{Z}}\int_{{\mathbb{R}}\setminus((-d_{1},d_{1})}\left|\mathbb{L}_{n_{2}}v_{n_{2}}(x)\right|^{2}dx=\mathcal{Z}_{u,1}^{2}.$
Using the proof of Lemma 6.5 in [6], we obtain that
$\mathcal{Z}_{u,2}^{2}\leq\mathcal{Z}_{u,1}^{2}+\sum_{n\in\mathbb{Z}^{2},n_{1}\neq 0}C_{n_{2}}^{2}\int_{(-d_{1},d_{1})^{2}}|v_{n_{2}}(x)v_{n_{2}}(z)|I_{n}(x,z)dxdz$
where
$I_{n}(x,z)\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{{\mathbb{R}}\setminus((-d_{1},d_{1})\cup(-d_{1},d_{1})+2d_{1}n_{1})}e^{-a_{n_{2}}|y-x|}e^{-a_{n_{2}}|2d_{1}n_{1}+z-y|}dy.$
We conclude the proof following the steps of the proof of Lemma 6.5 in [6] and using the above computations concerning the bound $\mathcal{Z}_{u,1}$. ∎
5 Stability Analysis
In this section, we aim at investigating the stability of traveling wave solutions to (2), of which the existence will be constructively established in Section 6. In particular, we assume that we know the existence of a solution $\tilde{u}\in\mathcal{H}$ to (2) such that
$\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0},$ (36)
where $\bar{u}$ is an approximate solution constructed as in (20) and $r_{0}$ is explicit. In order to study stability, we heavily rely on the framework developed in [18]. In fact, the authors of [18] investigated the orbital stability of traveling wave solutions in the 1D version of (2). One can easily show that the analysis of Sections 3 and 4 of [18] generalizes to the 2D version (2). We recall some notations for completeness.
Let $\Omega\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}{\mathbb{R}}\times(-d_{2},d_{2})$ be our spatial domain. Then, let $X\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}H^{2}(\Omega)\times L^{2}(\Omega)$ and let $X^{*}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}H^{2}(\Omega)\times L^{2}(\Omega)$ be its dual. Moreover, let us define the functional $E:X\to{\mathbb{R}}$ as
$E\begin{pmatrix}u\\ v\end{pmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\Omega}\frac{1}{2}v(x)^{2}+\frac{1}{2}(\Delta u(x))^{2}+e^{u(x)}-u(x)-1dx$ (37)
Then, (1) can be written in the form
$\frac{d}{dt}U=JE^{\prime}(U)$ (38)
where $U=\begin{pmatrix}u\\ v\end{pmatrix}$, $J\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}v\\ -u\end{pmatrix}$ and $E^{\prime}\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\Delta^{2}u+e^{u}-1\\ v\end{pmatrix}\in X^{*}$. Note that
$\tilde{U}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{pmatrix}\tilde{u}\\ c\partial_{x_{1}}\tilde{u}\end{pmatrix}$ (39)
is a solution to (38). Now, let $T(s):X\to X$ be the translation by $s\in\mathbb{R}$ in the direction $x_{1}$, that is $T(s)U(x)=U(x_{1}+s,x_{2})$. In particular, $T(s)\tilde{U}$ is a stationary solution to (38) for all $s\in\mathbb{R}$ by translation invariance.
We are now in a position to introduce the notation of orbital stability used in [18].
Definition 5.1.
Let $\tilde{u}$ be a traveling wave solution to (2) in $\mathcal{H}$ and let $\tilde{U}$ be defined as in (39). Then $\tilde{u}$ is said to be orbitally stable if the solution $U$ to (38) with initial condition $U(0)=U_{0}$ exists for all positive times and if for all $\epsilon>0$, there exists $\delta>0$ such that if $\|U_{0}-\tilde{U}\|_{X}<\delta$, then $\inf_{s\in{\mathbb{R}}}\|U(t)-T(s)\tilde{U}\|_{X}<\epsilon.$
$\tilde{u}$ is called orbitally unstable if it is not orbitally stable.
[…]
Lemma 5.1.
Let $\tilde{w}\in\mathcal{H}$ be defined as
$\tilde{w}=-2cD\mathbb{F}(\tilde{u})^{-1}\partial_{x_{1}}^{2}\tilde{u}$. (40)
Moreover, define the constant $\theta$ as follows
$\theta\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\Omega}(\tilde{u}+2c\tilde{w})\partial_{x_{1}}^{2}\tilde{u}$. (41)
If $n(D\mathbb{F}(\tilde{u}))=0$, or if $\theta>0$ and $n(D\mathbb{F}(\tilde{u}))=1$, then $\tilde{u}$ is orbitally stable. If $\theta>0$ and $)n(D\mathbb{F}(\tilde{u}$)) is even, or if $\theta<0$ and $)n(D\mathbb{F}(\tilde{u}$)) is odd, then $\tilde{u}$ is orbitally unstable.
Remark 5.2.
Using the notations of [18], $\tilde{w}$ in Lemma 5.1 corresponds to $v_{c}$ in Section 4.3 of [18]. Moreover, $\theta$ corresponds to $d^{\prime\prime}(c)$, after using integrations by parts (cf. Section 4.3 of [18]).
[…]
$\sigma_{ess}(D\mathbb{F}(\tilde{u})=\{l_{n_{2}}(\xi),~\xi\in{\mathbb{R}},n_{2}\in\mathbb{N}_{0}\}\subset\left[1-\frac{c^{4}}{4},\infty\right).$
Then, using Lemma 2.2 of [4], we have the following disjoint decomposition of the spectrum
$).\displaystyle\sigma(D\mathbb{F}(\tilde{u}))=\sigma_{ess}(D\mathbb{F}(\tilde{u}))\cup Eig(D\mathbb{F}(\tilde{u}$)). (42)
[…]
Lemma 5.3.
Let $\lambda_{min}$ be defined as
$\lambda_{min}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}-e^{\kappa_{2}r_{0}}\|e^{\bar{U}}\|_{1}-\frac{c^{4}}{4}.$ (43)
Then $(-\infty,1-\frac{c^{4}}{4})\cap Eig(D\mathbb{F}(\tilde{u}))\subset[\lambda_{min},1-\frac{c^{4}}{4})$.
Proof.
Let $)\lambda\in Eig(D\mathbb{F}(\tilde{u}$)). Then, there exists $u\in L^{2}(\Omega)$, $u\neq 0$, such that
$\mathbb{L}u+(e^{\tilde{u}}-1)u=\lambda u.$
This implies that
$|1-\frac{c^{4}}{4}-\lambda|\|u\|_{2}\leq\|\mathbb{L}u-\lambda u\|_{2}\leq\|e^{\tilde{u}}u-u\|_{2}\leq\|e^{\tilde{u}}u\|_{2}+\|u\|_{2}.$
Now, we have
$\|e^{\tilde{u}}u\|_{2}=\|e^{\tilde{u}-\bar{u}}e^{\bar{u}}u\|_{2}$ $\leq\|e^{\tilde{u}-\bar{u}}\|_{\infty}\|e^{\bar{u}}\|_{\infty}\|u\|_{2}$
$\leq e^{\kappa_{2}\|\tilde{u}-\bar{u}\|_{\mathcal{H}}}\|e^{\bar{U}}\|_{1}\|u\|_{2}$
$\leq e^{\kappa_{2}r_{0}}\|e^{\bar{U}}\|_{1}\|u\|_{2}.$
This concludes the proof. ∎
[…]
Using the ideas of Section 5.1 in [4], we notice that $L^{2}$ can be decomposed as follows
$L^{2}=L^{2}_{e}\oplus L^{2}_{o},$
where
$L^{2}_{o}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{u\in L^{2},~u(x_{1},x_{2})=-u(-x_{1},x_{2})\text{ for all }x\in{\mathbb{R}}^{2}\right\}.$ (44)
Moreover, given $\tilde{u}\in\mathcal{H}$, we have that
$D\mathbb{F}(\tilde{u})v_{i}\in L^{2}_{i}$
for all $v_{i}\in L^{2}_{i}$ and all $i\in\{e,o\}$. This implies that if $u\in L^{2}$ satisfies
$D\mathbb{F}(\tilde{u})u=\lambda u$
and we decompose $u=u_{e}+u_{o}$, where $u_{i}\in L^{2}_{i}$, then
$D\mathbb{F}(\tilde{u})u_{i}=\lambda u_{i}$
[…]
$\mathcal{D}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}P^{-1}DF(\bar{U})P,$
for some invertible infinite matrix $P=P^{N}+\pi_{N}$, with $P^{N}=\pi^{N}P^{N}\pi^{N}$, whose inverse is given as $P^{-1}=\pi^{N}(P^{N})^{-1}\pi^{N}+\pi_{N}$. $(P^{N})^{-1}$ has to be understood as the inverse of $P^{N}:\pi^{N}\ell^{2}\to\pi^{N}\ell^{2}$. In practice, columns of $P$ are approximate eigenvectors of $DF(\bar{U})$, so that $\mathcal{D}$ is close to being diagonal. We define $S$ as the diagonal part of the matrix $\mathcal{D}$ and $R=\mathcal{D}-S$. Moreover, we define $\lambda_{k}$ as the diagonal entries of $S$, that is
$(SU)_{k}=\lambda_{k}U_{k}\text{ for all }k\in\mathbb{N}_{0}^{2}.$ (46)
[…]
Lemma 5.4.
Let $n_{2}\in\mathbb{N}_{0}$, and let us define $a_{n_{2}}(\mathcal{J})$, $b_{n_{2}}(\mathcal{J})$ and ${C}_{n_{2}}(\mathcal{J})$ as
$a_{n_{2}}(\mathcal{J})$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\sqrt{4(1+\tilde{n}_{2}^{2}c^{2}-\delta_{0})-c^{4}}}{2\sqrt{c^{2}-2\tilde{n}_{2}^{2}+\sqrt{(c^{2}-2\tilde{n}_{2}^{2})^{2}+4(1+c^{2}\tilde{n}_{2}^{2}-\delta_{0})-c^{4}}}},$ (47)
$b_{n_{2}}(\mathcal{J})$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2}\sqrt{c^{2}-2\tilde{n}_{2}^{2}+\sqrt{(c^{2}-2\tilde{n}_{2}^{2})^{2}+4(1+c^{2}\tilde{n}_{2}^{2}-\delta_{0})-c^{4}}},$ (48)
$C_{n_{2}}(\mathcal{J})$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2(a_{n_{2}}(\mathcal{J})^{2}+b_{n_{2}}(\mathcal{J})^{2})^{\frac{1}{2}}\pi\sqrt{4(1+\tilde{n}_{2}^{2}c^{2}-\delta_{0})-c^{4}}}.$ (49)
Moreover, given $\mu\in\mathcal{J}$, we define $f_{n_{2},\mu}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{F}^{-1}(\frac{1}{l_{n_{2}}-\mu})$. Then, we obtain that
$\sup_{\mu\in\mathcal{J}}|f_{n_{2},\mu}(x)|\leq C_{n_{2}}(\mathcal{J})e^{-a_{n_{2}}(\mathcal{J})|x|},~~\text{ for all }x\in{\mathbb{R}}.$ (50)
Proof.
The proof is very similar to the one of Lemma 4.6. Indeed, we compute $f_{n_{2},\mu}$ as in the proof of Lemma 4.6, and obtain that
$|f_{n_{2},\mu}(x)|\leq C_{n_{2}}(\mu)e^{-a_{n_{2}}(\mu)|x|},~~\text{ for all }x\in{\mathbb{R}}$
where
$a_{n_{2}}(\mu)$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\sqrt{4(1+\tilde{n}_{2}^{2}c^{2}-\mu)-c^{4}}}{2\sqrt{c^{2}-2\tilde{n}_{2}^{2}+\sqrt{(c^{2}-2\tilde{n}_{2}^{2})^{2}+4(1+c^{2}\tilde{n}_{2}^{2}-\mu)-c^{4}}}},$
$b_{n_{2}}(\mu)$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2}\sqrt{c^{2}-2\tilde{n}_{2}^{2}+\sqrt{(c^{2}-2\tilde{n}_{2}^{2})^{2}+4(1+c^{2}\tilde{n}_{2}^{2}-\mu)-c^{4}}},$
$C_{n_{2}}(\mu)$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2(a_{n_{2}}(\mu)^{2}+b_{n_{2}}(\mu)^{2})^{\frac{1}{2}}\pi\sqrt{4(1+\tilde{n}_{2}^{2}c^{2}-\mu)-c^{4}}}.$
Then, one can prove that both $a_{n_{2}}$ and $b_{n_{2}}$ are decreasing functions of $\mu$, leading to the desired result. ∎
[…]
Lemma 5.5.
Let $t>-\lambda_{min}$ and let $\mathcal{J}=[\lambda_{min},\delta_{0}]$. Moreover, let us define $E=(E_{n})_{n\in\mathbb{N}_{0}^{2}}$ as follows:
$E_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{(C_{n_{2}}(\mathcal{J}))^{2}a_{n_{2}}(\mathcal{J})(-1)^{n_{1}}(1-e^{-4a_{n_{2}}(\mathcal{J})d_{1}})}{d_{1}(4a_{n_{2}}(\mathcal{J})^{2}+(2\pi\tilde{n}_{1})^{2})}.$ (51)
Furthermore, define $a(\mathcal{J})$ and $C(d_{1})$ as
$a(\mathcal{J})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\inf_{n_{2}\in\mathbb{N}_{0}}a_{n_{2}}(\mathcal{J})~\text{ and }~C(d_{1})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4d_{1}+\frac{4e^{-a(\mathcal{J})d_{1}}}{a(\mathcal{J})(1-e^{-\frac{3}{2}a(\mathcal{J})d_{1}})}+\frac{2}{a(\mathcal{J})(1-e^{-2a(\mathcal{J})d_{1}})}.$
Now, let us introduce various constants satisfying the following
$\mathcal{Z}_{u,1}$ $\geq 4d_{1}(\bar{V},\bar{V}*E),$
$\mathcal{Z}_{u,2}$ $\geq 4d_{1}(\bar{V},\bar{V}*,E)_{2}+2C(d_{1})e^{-2a(\mathcal{J})d_{1}}(\bar{V},\bar{V}*E)_{2}$
$\mathcal{Z}_{u,3}$ $\geq\|\pi^{N}(S+tI)^{-1}P^{-1}(L-\delta_{0}I)\|_{2}\mathcal{Z}_{u,2},$
$\mathcal{C}_{1}$ $\geq\frac{\|e^{\bar{U}}\|_{1}}{1-\delta_{0}-\frac{c^{4}}{4}}\frac{e^{\kappa_{2}r_{0}}-1}{r_{0}},$
$\mathcal{C}_{2}$ $\geq\|e^{\bar{U}}\|_{1}\frac{\|\pi^{N}(S+tI)^{-1}P^{-1}(L-\delta_{0}I)\|_{2}}{1-\delta_{0}-\frac{c^{4}}{4}}\frac{e^{\kappa_{2}r_{0}}-1}{r_{0}},$
$Z_{1,1}$ $\geq\frac{1}{1-\delta_{0}-\frac{c^{4}}{4}}\|\pi_{N}(L-\delta_{0}I)^{-1}R\pi^{N}\|_{2},~~~~Z_{1,2}\geq\frac{1}{1-\delta_{0}-\frac{c^{4}}{4}}\|\pi_{N}(L-\delta_{0}I)^{-1}R\pi_{N}\|_{2},$
$Z_{1,3}$ $\geq\|\pi^{N}(S+tI)^{-1}R\pi^{N}\|_{2},~~~~Z_{1,4}\geq\|\pi^{N}(S+tI)^{-1}R\pi_{N}\|_{2}.$
If
$\mathcal{C}_{1}r_{0}<1,$ (52)
we define $\beta_{1}$ as $\beta_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{u,1}+\mathcal{C}_{1}r_{0}}{1-\mathcal{C}_{1}r_{0}}.$ If in addition
$1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}>0,$ (53)
then we define
$\beta_{2}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{Z_{1,1}+(\mathcal{Z}_{u,2}+(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0})\|P^{N}\|_{2}}{1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}},$
$\epsilon_{n}^{(q)}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|\lambda_{n}+t|\left(Z_{1,3}+Z_{1,4}\frac{Z_{1,1}+\mathcal{Z}_{u,2}\|P^{N}\|_{2}}{1-Z_{1,2}-\mathcal{Z}_{u,2}}+\mathcal{Z}_{u,3}\left(\|P^{N}\|_{2}+\frac{Z_{1,1}+\mathcal{Z}_{u,2}\|P^{N}\|_{2}}{1-Z_{1,2}-\mathcal{Z}_{u,2}}\right)\right),$
$\epsilon_{n}^{(\infty)}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|\lambda_{n}+t|\left(Z_{1,3}+Z_{1,4}\beta_{2}+\left(\mathcal{Z}_{u,3}+\mathcal{C}_{2}r_{0}(1+\beta_{1}^{2})^{\frac{1}{2}}\right)\left(\|P^{N}\|_{2}+\beta_{2}\right)\right),$
$\epsilon_{n}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\max\{\epsilon_{n}^{(q)},\epsilon_{n}^{(\infty)}\}$
[…]
Proof.
The proof of the lemma follows the ones of Lemmas 4.1 and 4.2 in [4]. Now, we comment on the computation of the bounds. Following Lemma 4.1 in [4], we want to compute $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2}$ such that
$\mathcal{Z}_{u,1}$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\Omega\setminus\Omega_{0}}\left(\mathbb{L}-\mu I\right)^{-1}D\mathbb{G}(\bar{u})\right\|_{2},$
$\mathcal{Z}_{u,2}$ $\geq\sup_{\mu\in\mathcal{J}}\|\mathbb{1}_{\Omega_{0}}\left(\Gamma^{\dagger}\left((L-\mu I)^{-1}\right)-(\mathbb{L}-\mu I)^{-1}\right)D\mathbb{G}(\bar{u})\|_{2}.$
Now, noticing that $\left(\mathbb{L}_{n_{2}}-\mu I\right)^{-1}$ is the convolution operator associated to the function $f_{n_{2},\mu}$ defined in Lemma 5.4, we combine Lemma 4.7 and Lemma 5.4 to obtain that
$2d_{1}(\bar{V},\bar{V}*E)$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\Omega\setminus\Omega_{0}}\left(\mathbb{L}-\mu I\right)^{-1}D\mathbb{G}(\bar{u})\right\|_{2}$
$2d_{1}(\bar{V},\bar{V}*E)_{2}+C(d_{1})e^{-2ad_{1}}(\bar{V},\bar{V}*E)_{2}$ $\geq\sup_{\mu\in\mathcal{J}}\|\mathbb{1}_{\Omega_{0}}\left(\Gamma^{\dagger}\left((L-\mu I)^{-1}\right)-(\mathbb{L}-\mu I)^{-1}\right)D\mathbb{G}(\bar{u})\|_{2}.$
In particular, this validates our choice for $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2}$. In fact, as explained in the proof of Theorem 5.2 in [4], the extra factor $2$ in the bounds $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2}$ allows them to satisfy the inequalities for $\mathcal{Z}_{u,1}^{(q)},\mathcal{Z}_{u,2}^{(q)}$ in Lemma 4.2 in [4]. This means that we can choose $\mathcal{Z}_{u,1}^{(q)}=\mathcal{Z}_{u,1}$ and $\mathcal{Z}_{u,2}^{(q)}=\mathcal{Z}_{u,2}$ in the aforementioned lemma. For the bound $\mathcal{Z}_{u,3}$, we need
$\mathcal{Z}_{u,3}$ $\geq\sup_{\mu\in\mathcal{J}}\|\pi^{N}(S+tI)^{-1}P^{-1}(L-\mu I)\|_{2}\|\mathbb{1}_{\Omega_{0}}\left(\Gamma^{\dagger}\left((L-\mu I)^{-1}\right)-(\mathbb{L}-\mu I)^{-1}\right)D\mathbb{G}(\bar{u})\|_{2}.$
Let $\mu\in\mathcal{J}$, then we have
$\|\pi^{N}(S+tI)^{-1}P^{-1}(L-\mu I)\|_{2}\leq\|\pi^{N}(S+tI)^{-1}P^{-1}(L-\delta_{0}I)\|_{2}\|(L-\delta_{0}I)^{-1}(L-\mu I)\|_{2}.$
But then, we have
$\|(L-\delta_{0}I)^{-1}(L-\mu I)\|_{2}=\sup_{n\in\mathbb{N}_{0}^{2}}\frac{l(\tilde{n})-\mu}{l(\tilde{n})-\delta_{0}}=\sup_{n\in\mathbb{N}_{0}^{2}}1+\frac{\delta_{0}-\mu}{l(\tilde{n})-\delta_{0}}=\frac{1-\frac{c^{4}}{4}-\mu}{1-\frac{c^{4}}{4}-\delta_{0}}$
using Lemma 7.1. Then, recalling $C_{n_{2}}(\mu)$ from the proof of Lemma 5.4, one can prove that $\mu\mapsto C_{n_{2}}(\mu)\frac{1-\frac{c^{4}}{4}-\mu}{1-\frac{c^{4}}{4}-\delta_{0}}$ is increasing in $\mu$ and consequently the maximum is reached at $\mu=\delta_{0}$ when considering $\mu\in\mathcal{J}$. This concludes the obtained inequality for $\mathcal{Z}_{u,3}$. Now, the computations of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ follow directly from the proof of Lemma 4.3. In particular, we also use that $\|(\mathbb{L}-\mu I)^{-1}\|_{2}\leq\frac{1}{1-\frac{c^{4}}{4}-\delta_{0}}$ for all $\mu\in\mathcal{J}$. The rest of the proof follows from [4]. ∎
[…]
5.2 Enclosure of $\theta$
In this section, we provide a computer-assisted strategy for enclosing the value of $\theta$ defined in Lemma 5.1. In fact, we derive the following result.
Lemma 5.6.
Let $r_{0}>0$ and let $\tilde{u}\in\mathcal{H}$ be a solution to (2) such that $\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}$, where $\bar{u}\in\mathcal{H}$ is constructed as in (20). Let $\overline{W}\in\pi^{N}\ell^{2}$ such that $\overline{w}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(\overline{W})\in\mathcal{H}$. Let $\epsilon_{0},\epsilon>0$ be defined as
$\epsilon_{0}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\kappa_{1}r_{0}+4c^{2}\|D\mathbb{F}(\tilde{u})^{-1}\|_{2}\left(\|\partial_{x_{1}}\bar{u}+\frac{1}{2c}D\mathbb{F}(\bar{u})\overline{w}\|_{2}+\frac{r_{0}}{2-c^{2}}+\frac{1}{2c}(1-e^{\kappa_{2}r_{0}})\|e^{\bar{U}}\overline{W}\|_{1}\right),$
$\epsilon$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\|\bar{u}+2c\overline{w}\|_{2}\|\partial_{x_{1}}^{2}\mathbb{L}^{-1}\|_{2}r_{0}+\epsilon_{0}\left(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\frac{r_{0}}{2-c^{2}}\right).$ (54)
Then, we obtain that
$\theta\in[\theta_{0}-\epsilon,\theta_{0}+\epsilon],$ (55)
where
$\theta_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}|{\Omega_{0}}|\left((\bar{U}+2c\overline{W}),\partial_{x_{1}}^{2}\bar{U}\right)_{2},~~\text{ and where }(\partial_{x_{1}}^{2}\bar{U})_{n}=-\frac{n_{1}^{2}\pi^{2}}{d_{1}^{2}}\bar{U}_{n}\text{ for all }(n_{1},n_{2})\in\mathbb{N}_{0}^{2}.$
Proof.
Recalling the definition of $\theta$ in (41), we have
$\theta=\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}\bar{u}+\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})+\int_{\Omega}(\tilde{u}-\overline{u}+2c\tilde{w}-2c\overline{w})\partial_{x_{1}}^{2}\tilde{u}$,
where $\overline{w}\in\mathcal{H}$ is an approximation of $\tilde{w}$. In particular, using Cauchy-Schwarz inequality, we have
$\left|\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})\right|$ $\leq\|\bar{u}+2c\overline{w}\|_{2}\|\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})\|_{2}$
$\leq\|\bar{u}+2c\overline{w}\|_{2}\|\partial_{x_{1}}^{2}\mathbb{L}^{-1}\|_{2}\|\mathbb{L}(\tilde{u}-\bar{u})\|_{2}$
$\leq\|\bar{u}+2c\overline{w}\|_{2}\|\partial_{x_{1}}^{2}\mathbb{L}^{-1}\|_{2}r_{0}$
using that $\|\mathbb{L}(\tilde{u}-\bar{u})\|_{2}=\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}.$ Moreover, using the Fourier transform, we have that
$\|\partial_{x_{1}}^{2}\mathbb{L}^{-1}\|_{2}\leq\sup_{\xi\in{\mathbb{R}}^{2}}\frac{(2\pi\xi_{1})^{2}}{l(\xi)}\leq\frac{1}{2-c^{2}}.$
On the other hand, we have
[…]
$\leq(\|\tilde{u}-\overline{u}\|_{2}+2c\|\tilde{w}-\overline{w}\|_{2})(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\|\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})\|_{2})$
$\leq(\kappa_{1}r_{0}+2c\|\tilde{w}-\overline{w}\|_{2})(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\frac{r_{0}}{2-c^{2}}).$
Consequently, it remains to compute $\|\tilde{w}-\overline{w}\|_{2}$. By definition of $\tilde{w}$ in (40), we have
$\|\tilde{w}-\overline{w}\|_{2}=\|-2cD\mathbb{F}(\tilde{u})\partial_{x_{1}}^{2}\tilde{u}-\overline{w}\|_{2}\leq 2c\|D\mathbb{F}(\tilde{u})^{-1}\|_{2}\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
Now, we have
$\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}\leq\|\partial_{x_{1}}\bar{u}+\frac{1}{2c}D\mathbb{F}(\bar{u})\overline{w}\|_{2}+\|\partial_{x_{1}}^{2}(\bar{u}-\tilde{u})\|_{2}+\frac{1}{2c}\|D\mathbb{F}(\bar{u})\overline{w}-D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
Moreover, we obtain that
$\|D\mathbb{F}(\bar{u})\overline{w}-D\mathbb{F}(\tilde{u})\overline{w}\|_{2}=\|(e^{\bar{u}}-e^{\tilde{u}})\overline{w}\|_{2}\leq\|1-e^{\tilde{u}-\bar{u}}\|_{2}\|e^{\bar{u}}\overline{w}\|_{\infty}$
$\leq(1-e^{\kappa_{2}r_{0}})\|e^{\bar{U}}\overline{W}\|_{1}.$
Finally, using that $\bar{u}=\gamma^{\dagger}(\bar{U})$ and $\overline{w}=\gamma^{\dagger}(\overline{W})$, we use Parseval’s identity to get
$\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}\bar{u}=|{\Omega_{0}}|\left((\bar{U}+2c\overline{W}),\partial_{x_{1}}^{2}\bar{U}\right)_{2},$
where $(\partial_{x_{1}}^{2}\bar{U})_{n}=-\frac{n_{1}^{2}\pi^{2}}{d_{1}^{2}}\bar{U}_{n}$ for all $(n_{1},n_{2})\in\mathbb{N}_{0}^{2}$. This concludes the proof. ∎
Using the previous result, we obtain a computer-assisted strategy for the enclosure of the value of $\theta$. Indeed, the constant $\epsilon$ depends on finite-dimensional computations, which can be handled using rigorous numerics (cf. [31]). In particular, this allows us to conclude about the stability of $\tilde{u}$.
6 Results
In this section, we present the main results of the paper, which consist of rigorous, computer-assisted bounds on numerical approximations to (2). These approximate solutions were obtained using Newton’s method, initialized with carefully constructed guesses based on approximations in the periodic setting described in [32]. To extend the initial guesses in the periodic case to the infinite setting, we project them into the kernel of matrix $\mathcal{T}$, as described in Section 3.1.
By combining these new approximations with the estimates established in the previous section, we obtain a validated proof of existence for solutions to the suspension bridge equation within explicitly controlled error bounds. This leads to the following theorem statement:
Theorem 6.1.
For $c=1.2$, $d=(\frac{\pi}{0.06},\frac{\pi}{0.24})$, $N_{0}=(150,80)$, $N=(40,40)$, there exists a solution $\tilde{u}\in\mathcal{H}$ to (2) on $\mathbb{R}\times(-d_{2},d_{2})$ with Neumann boundary conditions in the $x_{2}$-direction such that
$\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq 9.5\cdot 10^{-8},$
where $\bar{u}$ is a numerically computed approximation visualized in Figure 1. Moreover, $D\mathbb{F}(\tilde{u}):L^{2}(\Omega)\to L^{2}(\Omega)$ possesses exactly one negative eigenvalue and $\theta>0$, where $\theta$ is given in (41). Finally, $\tilde{u}$ is orbitally stable.
Proof.
Recall that $\nu$ and $N^{\mathrm{FFT}}$ are defined in Section 3.2. We choose $\nu=(1.1,1.1)$ and $N^{\mathrm{FFT}}=(512,512)$. Finitely many cosine coefficients of the approximate solution $\bar{u}$, visualized in Figure 1, can be found in the code in [31]. We report the values for the bounds in floating point numbers displaying four decimals. We have
$\mathcal{Y}_{0}=6.0005\cdot 10^{-8},\quad\mathcal{Z}_{1}=3.6608\cdot 10^{-1},\quad\mathcal{Z}_{2}=22.7926.$
Letting $\tilde{u}$ denote the zero of $\mathbb{F}$, we apply Theorem 3.1 to obtain the error bound
$\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq 9.4657\cdot 10^{-8}.$
[…]
$0=\partial_{x_{1}}\mathbb{F}(\tilde{u})=D\mathbb{F}(\tilde{u})\partial_{x_{1}}\tilde{u}$. (56)
[…]
The code used to obtain the above result, as well as the results presented later in this section, can be found in [31]. The computations were carried out in Julia (v1.10.4) on an Apple M1 Pro CPU with $32$ GB RAM. We heavily relied on interval arithmetic using the packages IntervalArithmetic.jl ([26]) and RadiiPolynomial.jl ([17]) to ensure that the computations are rigorous.
The method developed in this paper is versatile and can also be applied to rigorously prove existence of solutions on infinite strips for other parameter values (e.g. different $c$, $d_{1}$, $d_{2}$) and with alternative spatial patterns. Two illustrative examples are presented in Figure 2. In contrast to the approximation shown in Figure 1, which contains a single prominent peak at the bottom, the approximations in Figure 2 show an alternative spatial pattern characterized by two dominant peaks at the bottom. In Figure 2(a), the parameter $c$ is slightly increased relative to the value used in Theorem 6.1, while in Figure 2(b), it is decreased. The plots indicate that reducing the value of $c$ leads to an increase in the amplitude of the resulting wave patterns.
Figure 2: Visualization of approximate two-peak solutions $\bar{u}$ to (2). The approximations are truncated to a finite domain in these plots.
The existence theorems and corresponding proofs, with stability analysis, of the two alternative patterns are given below.
Theorem 6.2.
For $c=1.3$, $d=(\frac{\pi}{0.06},\frac{\pi}{0.2})$, $N_{0}=(100,60)$, $N=(60,60)$, there exists a solution $\tilde{u}\in\mathcal{H}$ to (2) on $\mathbb{R}\times(-d_{2},d_{2})$ with Neumann boundary conditions in the $x_{2}$-direction such that
$\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq 3.1\cdot 10^{-6},$
where $\bar{u}$ is a numerically computed approximation visualized in Figure 2(a). Moreover, $D\mathbb{F}(\tilde{u}):L^{2}(\Omega)\to L^{2}(\Omega)$ possesses exactly two negative eigenvalues and $\theta>0$, where $\theta$ is given in (41). Finally, $\tilde{u}$ is orbitally unstable.
Proof.
We choose $\nu=(1.1,1.1)$ and $N^{\mathrm{FFT}}=(512,512)$. We have
$\mathcal{Y}_{0}=2.4835\cdot 10^{-6},\quad\mathcal{Z}_{1}=1.792\cdot 10^{-1},\quad\mathcal{Z}_{2}=40.3926.$
Letting $\tilde{u}$ denote the zero of $\mathbb{F}$, we apply Theorem 3.1 to obtain the error bound
$\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq 3.0256\cdot 10^{-6}.$
The stability part follows similarly as the one presented in the proof of Theorem 6.1. In this case, we provide the existence of exactly two negative eigenvalues by proving that two Gershgorin intervals of Lemma 5.5 are in the negative part of $\mathbb{R}$}, disjoint from on another, and disjoint from the rest of the intervals. ∎
Theorem 6.3.
For $c=0.8$, $d=(\frac{\pi}{0.1},\frac{\pi}{0.2})$, $N_{0}=(80,60)$, $N=(80,60)$, there exists a solution $\tilde{u}\in\mathcal{H}$ to (2) on $\mathbb{R}\times(-d_{2},d_{2})$ with Neumann boundary conditions in the $x_{2}$-direction such that
$\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq 4.0\cdot 10^{-5},$
where $\bar{u}$ is a numerically computed approximation visualized in Figure 2(b). Moreover, $D\mathbb{F}(\tilde{u}):L^{2}(\Omega)\to L^{2}(\Omega)$ possesses exactly two negative eigenvalues and $\theta>0$, where $\theta$ is given in (41). Finally, $\tilde{u}$ is orbitally unstable.
Proof.
We choose $\nu=(1.1,1.1)$ and $N^{\mathrm{FFT}}=(512,512)$. We have
$\mathcal{Y}_{0}=3.3557\cdot 10^{-5},\quad\mathcal{Z}_{1}=1.4034\cdot 10^{-1},\quad\mathcal{Z}_{2}=65.7018.$
Letting $\tilde{u}$ denote the zero of $\mathbb{F}$, we apply Theorem 3.1 to obtain the error bound
$\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq 3.9094\cdot 10^{-5}.$
For the stability part, we obtain $\lambda_{min}\approx-15$, using the notation of Lemma 5.3. Then, we choose $t>\lambda_{min}$ in Lemma 5.5 and compute the bounds of the lemma to compute the Gershgorin intervals $\overline{B_{\epsilon_{j}}(\lambda_{j})}$. We obtained intersecting intervals and could not conclude about the localization of negative eigenvalues. However, we obtained that the union of all the intervals was contained in $(-0.4,\infty)$. This implies that we can improve our choice for $\lambda_{min}$ and choose $\lambda_{min}=-0.45$. Choosing $t=0.5$ in Lemma 5.3 and computing the bounds again, we get tighter radiae for the Gershgoring intervals and obtain disjoint intervals on the negative part of $\mathbb{R}$}.
The stability part then follows similarly as the one presented in the proof of Theorem 6.1. In this case, we provide the existence of exactly two negative eigenvalues by proving that two Gershgorin intervals of Lemma 5.5 are in the negative part of $\mathbb{R}$}, disjoint from one another, and disjoint from the rest of the intervals. ∎
The computation time to obtain the result in Theorem 6.1 was $892$ seconds, including the stability proof. The computation times corresponding to the results in Theorems 6.2 and 6.3 were $420$ and $586$ seconds, respectively. The increased computation time associated with our first theorem arises from selecting a larger value for $N_{0}$, which in turn yields a tighter error bound.
7 Appendix
In the next lemma we provide some technical results about $l_{n}$.
Lemma 7.1.
We have
$\min_{\xi_{1}\in{\mathbb{R}}}l_{n_{2}}(\xi_{1})=\begin{cases}c^{2}\tilde{n}_{2}^{2}+1-\frac{c^{4}}{4}&\text{ if }0\leq n_{2}\leq\frac{d_{2}c}{\pi\sqrt{2}}\\ \tilde{n}_{2}^{4}+1&\text{ otherwise.}\end{cases}$ (57)
Furthermore, given $n_{2}\geq\frac{d_{2}c}{\pi}$, we have
$l_{n_{2}}(\xi_{1})\geq\frac{(2\pi)^{4}}{2}(\tilde{n}_{2}^{2}+\xi_{1}^{2})^{2}$
for all $\xi\in{\mathbb{R}}$. In particular, given $N\in\mathbb{N}$ such that $N>\frac{d_{2}c}{\pi}$, this implies that
$\sum_{n_{2}=N+1}^{\infty}\left\|\frac{1}{l_{n_{2}}}\right\|_{2}^{2}\leq\frac{5(2d_{2})^{7}}{48(2\pi)^{7}N^{6}}.$ (58)
Proof.
Denoting $x=(2\pi\xi_{1})^{2}$ and $\alpha=(2\pi\xi_{2})^{2}$, we have
$l(\xi_{1},\xi_{2})=(x+\alpha)^{2}-c^{2}x+1\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}p_{\alpha}(x).$
Studying the variations of $p_{\alpha}$, we have that $p_{\alpha}$ has a global minimum at $x=\frac{c^{2}}{2}-\alpha$. In particular, this implies that
$l(\xi_{1},\xi_{2})\geq\begin{cases}(2\pi\xi_{2})^{2}c^{2}+1-\frac{c^{4}}{4}&\text{ if }\frac{c^{2}}{2}\geq(2\pi\xi_{2})^{2}\\ (2\pi\xi_{2})^{4}+1&\text{ otherwise}.\end{cases}$
Choosing $\xi_{2}=\frac{n_{2}}{2d_{2}}$, we obtain the desired result. To prove the remaining statement, notice that
$l_{n_{2}}(\xi_{1})-\frac{1}{2}((2\pi\tilde{n}_{2})^{2}+(2\pi\xi_{1})^{2})^{2}$ $=\frac{1}{2}((2\pi\tilde{n}_{2})^{2}+x)^{2}-c^{2}x+1$
$=\frac{1}{2}((2\pi\tilde{n}_{2})^{4}+2(2\pi\tilde{n}_{2})^{2}x+x^{2})-c^{2}x+1$
where $x=(2\pi\xi_{1})^{2}$. Consequently, we have that $l_{n_{2}}(\xi_{1})-\frac{1}{2}((2\pi\tilde{n}_{2})^{2}+(2\pi\xi_{1})^{2})^{2}\geq 0$ if $2\pi\tilde{n}_{2}\geq c$. Now, we have
$\left\|\frac{1}{l_{n_{2}}}\right\|_{2}^{2}\leq\frac{4}{(2\pi)^{8}}\int_{{\mathbb{R}}}\frac{1}{(\tilde{n}_{2}^{2}+\xi_{1}^{2})^{4}}d\xi_{1}=\frac{20\pi}{16(2\pi)^{8}\tilde{n}_{2}^{7}}=\frac{5}{8(2\pi\tilde{n}_{2})^{7}}.$
Moreover, we get
$\sum_{n_{2}=N+1}^{\infty}\frac{1}{\tilde{n}_{2}^{7}}=(2d_{2})^{7}\sum_{n_{2}=N+1}^{\infty}\frac{1}{n_{2}^{7}}\leq(2d_{2})^{7}\int_{N}^{\infty}\frac{1}{x^{7}}dx=\frac{(2d_{2})^{7}}{6N^{6}}.$
Noticing that $2\pi\tilde{n}_{2}\geq c$ is equivalent to $n_{2}\geq\frac{d_{2}c}{\pi}$, we have
$2\sum_{n_{2}=N+1}^{\infty}\left\|\frac{1}{l_{n_{2}}}\right\|_{2}^{2}\leq\frac{(2d_{2})^{7}}{6N^{6}}\frac{5}{8(2\pi)^{7}}$
where $N=\lceil{\frac{d_{2}c}{\pi}}\rceil.$

### v3

@@
-Existence proofs of traveling wave solutions on an infinite strip for the suspension bridge equation and proof of orbital stability
+Existence and orbital stability proofs of traveling wave solutions on an infinite strip for the suspension bridge equation
 Abstract
@@
 1 Introduction
-The objective of this paper is to prove the existence of traveling wave solutions of the two-dimensional suspension bridge equation on an infinite strip $\Omega\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}{\mathbb{R}}\times(-d_{2},d_{2})$ for $d_{2}>0$.
+The objective of this paper is to prove the existence of traveling wave solutions of the two-dimensional suspension bridge equation on an infinite strip $\Omega\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}{\mathbb{R}}\times(-d_{2},d_{2})$ for $d_{2}>0$ and to prove orbital (in)stability.
 We let $v=v(t,X_{1},X_{2}):{\mathbb{R}}^{+}\times\Omega\to{\mathbb{R}}$ represent the deflection in the downward direction of the surface of the bridge and we let $\Delta^{2}=(\partial_{X_{1}}^{2}+\partial_{X_{2}}^{2})^{2}$.
@@
 $\partial_{x_{2}}u(\cdot,\pm d_{2})=\partial_{x_{2}}^{3}u(\cdot,\pm d_{2})=0.$ (3)
-We only consider values of $c$ in $(0,\sqrt{2})$, because if $c$ tends to zero, the amplitude of the waves grows closer to infinity, and for $c=0$, nontrivial solutions do not longer exist [23].
+We only consider values of $c$ in $(0,\sqrt{2})$, because if $c$ tends to zero, the amplitude of the waves grows to infinity, and for $c=0$, nontrivial solutions do not longer exist [23].
 Contrarily, if $c$ approaches $\sqrt{2}$, the solutions become oscillatory as the amplitudes of the waves tend to go to zero [16].
-Our goal is to establish the existence of a solution $\tilde{u}$ to (2) on $\Omega$.
+We want to establish the existence of a solution $\tilde{u}$ to (2) on $\Omega$.
 Numerical simulations of solutions to this equation, obtained via the Mountain Pass Algorithm, are discussed in [16].
@@
 Figure 1: Visualization of an approximate solution $\bar{u}$ to (2) with $c=1.2$. The approximation is truncated to a finite domain in this plot.
-Ultimately our goal is to construct an approximate solution $\bar{u}:\Omega\to{\mathbb{R}}$ to (2) and then prove the existence of a true solution to (2) in a vicinity of $\bar{u}$.
+In this paper, we construct an approximate solution $\bar{u}:\Omega\to{\mathbb{R}}$ to (2) and then prove the existence of a true solution to (2) in a vicinity of $\bar{u}$.
 An illustration of such an approximate solution $\bar{u}$, that we rigorously verify in Theorem 6.1, is depicted in Figure 1.
@@
 Specifically, we use a variant of this theorem called the radii polynomial theorem [10].
-To compensate for the fact that we are working on an unbounded domain, an additional bound that needs to be computed is introduced, on top of the other bounds mentioned in the radii polynomial theorem.
+To prove existence on a bounded domain using this theorem, we need to compute certain bounds.
+To compensate for the fact that we are working on an unbounded domain, an additional bound is introduced and computed in this paper.
 To guarantee the reliability of the computational components of the proof, we employ Julia interval arithmetic using the packages [15] and [26].
-The solutions obtained in the CAPs will also be analyzed for stability.
+The solutions obtained in our CAPs will also be analyzed for stability.
 This part of the work is based on the procedure outlined in [18], where the authors investigate orbital stability of one-dimensional traveling wave solutions to the suspension bridge equation.
@@
 and where $u_{n_{2}}:\mathbb{R}\to\mathbb{R}$ and $u_{n_{2}}$ is even. Corresponding to (5), we define $L^{2}_{e}$ as the following Hilbert space of $L^{2}$ even functions on the strip $\Omega$:
-$L^{2}_{e}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{u(x_{1},x_{2})=u_{0}(x_{1})+2\sum_{n_{2}\in\mathbb{N}}u_{n_{2}}(x_{1})\cos\left(2\pi\tilde{n}_{2}x_{2}\right),~u_{n_{2}}(x)=u_{n_{2}}(-x)\text{ and }\|u\|_{2}<\infty\right\}$
+$L^{2}_{e}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{u(x_{1},x_{2})=u_{0}(x_{1})+2\sum_{n_{2}\in\mathbb{N}}u_{n_{2}}(x_{1})\cos\left(2\pi\tilde{n}_{2}x_{2}\right),~u_{n_{2}}(x_{1})=u_{n_{2}}(-x_{1})\text{ and }\|u\|_{2}<\infty\right\}$
 where
@@
 where we recall that $\tilde{n}_{2}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{n_{2}}{2d_{2}}$.
-Now, it remains to prove that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is well defined in order to set-up (2) on $\mathcal{H}$. For that purpose, we use the following result:
+Now, it remains to prove that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is well defined in order to formulate (2) on $\mathcal{H}$. For that purpose, we use the following result:
 Lemma 2.1.
@@
 $\|u\|_{\infty}\leq\left(\left\|\frac{1}{l_{0}}\right\|_{L^{2}({\mathbb{R}})}^{2}+2\sum_{n_{2}\in\mathbb{N}}\left\|\frac{1}{l_{n_{2}}}\right\|_{L^{2}({\mathbb{R}})}^{2}\right)^{\frac{1}{2}}\|u\|_{\mathcal{H}}\text{ and }\|v\|_{2}\leq\max_{n_{2}\in\mathbb{N}_{0}}\sup_{\xi_{1}\in{\mathbb{R}}}\left|\frac{1}{l_{n_{2}}(\xi_{1})}\right|\|v\|_{\mathcal{H}}.$
-We conclude the proof using (57) in Lemma 7.1. ∎
+We conclude the proof using (57) from Lemma 8.1. ∎
 Remark 2.2.
@@
 In fact, this can be achieved using rigorous numerics (cf.
-LA: Reference code).
+[31]).
 To estimate the tail of the summation, we use (58) in Lemma 8.1.
-Using the above lemma, we have that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is a smooth operator. Accordingly, we investigate the solution of the following zero-finding problem
+Using the above lemma, we have that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is a smooth function. Accordingly, we investigate the solution of the following zero-finding problem
 $\mathbb{F}(u)=0~~\text{ for }u\in\mathcal{H},$ (8)
@@
 2.2 Fourier series representation
-As described in [6], existence proofs of localized solutions can be achieved thanks to approximate objects on Fourier coefficients.
-Following such an approach, we introduce notation corresponding to objects on Fourier coefficients, related to the ones introduced in the previous section.
-Let $d_{1}>0$ and define $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d_{1},d_{1})\times(-d_{2},d_{2})$.
-$\Omega_{0}\subset\Omega$ is the bounded domain on which our approximate solution is constructed.
+As described in [4], existence proofs of localized solutions can be achieved using approximations of the true solution in Fourier space.
+Following such an approach, we introduce notation corresponding to Fourier series, related to the ones introduced in the previous section.
+Let $d_{1}>0$ and let $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d_{1},d_{1})\times(-d_{2},d_{2})$ be the bounded domain on which our approximate solution is constructed.
 We let $\mathbb{N}_{0}^{2}=(\mathbb{N}\cup\{0\})^{2}$ and we define
@@
 Therefore, we define $\gamma:L^{2}_{e}\rightarrow\ell^{2}$ and $\gamma^{\dagger}:\ell^{2}\rightarrow L^{2}_{e}$ such that
-$\left((\gamma(u)\right)_{n}$ $=\frac{1}{|\Omega_{0}|}\int_{\Omega_{0}}u(x_{1},x_{2})e^{-2\pi i\tilde{n}\cdot x}dx\quad\text{ for }n\in\mathbb{N}_{0}^{2},$
+$\left(\gamma(u)\right)_{n}$ $=\frac{1}{|\Omega_{0}|}\int_{\Omega_{0}}u(x_{1},x_{2})e^{-2\pi i\tilde{n}\cdot x}dx\quad\text{ for }n\in\mathbb{N}_{0}^{2},$
 $\gamma^{\dagger}(U)(x)$ $=\mathbb{1}_{\Omega_{0}}(x)\sum_{n\in\mathbb{N}_{0}^{2}}\alpha_{n}U_{n}\cos\left(2\pi\tilde{n}_{1}x_{1}\right)\cos\left(2\pi\tilde{n}_{2}x_{2}\right),$
@@
 The map $\Gamma^{\dagger}$ will be useful when constructing operators on $\mathcal{B}(L^{2}_{e})$ thanks to operators on Fourier coefficients (that is, operators on $\mathcal{B}(\ell^{2})$).
-Furthermore, we introduce the norm $\|\cdot\|_{2,\mathcal{H}}$ to denote the operator norm for bounded linear operators on $L^{2}_{e}\rightarrow\mathcal{H}$.
+Furthermore, we introduce the norm $\|\cdot\|_{2,\mathcal{H}}$ to denote the operator norm for bounded linear operators $L^{2}_{e}\rightarrow\mathcal{H}$.
 Finally, given $u\in L^{\infty},U\in\ell^{1}$, we introduce notation to denote the linear multiplication operator $\mathbb{M}_{u}$ related to the function $u$ and the linear discrete convolution operator $\mathbb{M}_{U}$ related to the sequence $U$ as
@@
 Our approach is based on a Newton-Kantorovich theorem.
-More specifically, we construct a fixed point operator for which fixed points correspond to zeros of $\mathbb{F}$ (8).
+More specifically, we construct a fixed point operator for which fixed points correspond to zeros of $\mathbb{F}$.
 First, we construct an approximate solution $\bar{u}\in\mathcal{H}$ thanks to its Fourier coefficients representation.
 The technical construction of $\bar{u}$ is described in Section 3.1.
-Given $\bar{u}$, we then construct $\mathbb{A}:L^{2}_{e}\to\mathcal{H}$ as an approximate inverse for the Fréchet derivative $D\mathbb{F}(\bar{u})$.
+Given $\bar{u}$, we then construct an injective $\mathbb{A}:L^{2}_{e}\to\mathcal{H}$ as an approximate inverse for the Fréchet derivative $D\mathbb{F}(\bar{u})$.
 Then, our goal is to determine a radius $r>0$ such that $\mathbb{T}$, given as
@@
 is well defined on $\overline{B_{r}(\bar{u})}\to\overline{B_{r}(\bar{u})}$ and contracting (where $B_{r}(\bar{u})$ is the open ball of $\mathcal{H}$ centered at $\bar{u}$ and of radius $r$).
-The Banach fixed point theorem then provides the existence of a unique zero $\tilde{u}\in\mathcal{H}$ in $\overline{B_{r}(\bar{u})}.$
+The Banach fixed point theorem then provides the existence of a unique zero of $\tilde{u}\in\mathcal{H}$ of $\mathbb{F}$ in $\overline{B_{r}(\bar{u})}.$
 In order to verify the existence of such a radius $r$, we make use of the following result. Its proof is for instance given in [4].
@@
 That is, $\bar{U}$ has a vector representation of size $(N_{1}^{0}+1)(N_{2}^{0}+1).$
-Now, note that by construction $\bar{u}\in L^{2}_{e}$ but we do not necessarily have $\bar{u}\in\mathcal{H}.$ Indeed, $\bar{u}$ might not be smooth at $x_{1}=\pm d_{1}$.
-By periodicity of $\bar{u}$ on $\Omega_{0}$}, this is equivalent for asking smoothness only at $d_{1}$.
+Now, note that by construction $\bar{u}\in L^{2}_{e}$, but we do not necessarily have $\bar{u}\in\mathcal{H}.$ Indeed, $\bar{u}$ might not be smooth at $x_{1}=\pm d_{1}$.
+By periodicity of $\bar{u}$ on $\Omega_{0}$}, smoothness at $x_{1}=\pm d_{1}$ is equivalent to smoothness only at $d_{1}$.
 Now, in order to obtain the required regularity for $\mathcal{H}$, we need
 $\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0\text{ for all }x_{2}\in(-d_{2},d_{2})\text{ and all }k\in\{0,1,2,3\}.$
-Note that since $\bar{u}$ has a cosine-cosine representation, we readily have $\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0$ for all $x_{2}\in(-d_{2},d_{2})$ and all $k\in\{1,3\}$.
+Since $\bar{u}$ has a cosine-cosine representation, we readily have $\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0$ for all $x_{2}\in(-d_{2},d_{2})$ and all $k\in\{1,3\}$.
 Consequently, it remains to ensure that
 $\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0\text{ for all }x_{2}\in(-d_{2},d_{2})\text{ and all }k\in\{0,2\}.$ (19)
-Note that since $\bar{u}$ is represented thanks to $\bar{U}$ on $\Omega_{0}$}, (19) can be translated as equations on $\bar{U}$ :
+As $\bar{u}$ is represented by $\bar{U}$ on $\Omega_{0}$}, (19) can be translated to equations on $\bar{U}$ :
 $\sum_{n_{1}=-N^{0}_{1}}^{N^{0}_{1}}(-1)^{n_{1}}(2\pi\tilde{n}_{1})^{k}\bar{U}_{n_{1},n_{2}}=0\text{ for all }n_{2}\in-N^{0}_{2},\dots,N^{0}_{2}\text{ and all }k\in\{0,2\},$
-where we have used that $\bar{U}=\pi^{N}\bar{U}$. Note that this is equivalent to having $\bar{U}\in Ker(\mathcal{T})$, where
+where we have used that $\bar{U}=\pi^{N}\bar{U}$. Note that this is equivalent to having $\bar{U}\in\mathrm{Ker}(\mathcal{T})$, where
 $\mathcal{T}\bar{U}=\begin{pmatrix}\mathcal{T}_{0}\bar{U}\\ \mathcal{T}_{2}\bar{U}\end{pmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\begin{pmatrix}\left(\sum_{n_{1}=-N^{0}_{1}}^{N^{0}_{1}}(-1)^{n_{1}}\bar{U}_{n_{1},n_{2}}\right)_{n_{2}\in-N^{0}_{2},\dots,N^{0}_{2}}\\ \left(\sum_{n_{1}=-N^{0}_{1}}^{N^{0}_{1}}(-1)^{n_{1}}(2\pi\tilde{n}_{1})^{2}\bar{U}_{n_{1},n_{2}}\right)_{n_{2}\in-N^{0}_{2},\dots,N^{0}_{2}}\end{pmatrix}$.
-Consequently, $\bar{U}$ needs to be projected into the kernel of the matrix $\mathcal{T}$. Such an operation has been described in [6, 7] and is implemented in [31].
-Overall, we ensure that $\bar{U}\in Ker(\mathcal{T})$ and obtain that
+Consequently, $\bar{U}$ needs to be projected into the kernel of the matrix $\mathcal{T}$ of size $(N^{0}_{1}+1)(N^{0}_{2}+1)$.
+Such an operation has been described in [4, 5] and is implemented in [31].
+Overall, we ensure that $\bar{U}\in\mathrm{Ker}(\mathcal{T})$ and obtain that
 $\bar{u}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}\left(\bar{U}\right)\in\mathcal{H}~\text{ and }~\bar{U}=\pi^{N}\bar{U}.$ (20)
@@
 Following the procedure outlined in Section 3 of [32], we obtain the expression:
-$e^{\bar{U}}\in\begin{cases}{(e^{\bar{U}})}^{\mathrm{FFT}}_{n}+{C}[-\varepsilon_{n},\varepsilon_{n}]&\text{ if }0\leq n_{1}\leq N^{0}_{1},0\leq n_{2}\leq N^{0}_{2},\\ [-\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}},\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}}]&\text{ otherwise}\end{cases}$ (21)
+$\left(e^{\bar{U}}\right)_{n}\in\begin{cases}{(e^{\bar{U}})}^{\mathrm{FFT}}_{n}+{C}[-\varepsilon_{n},\varepsilon_{n}]&\text{ if }0\leq n_{1}\leq N^{0}_{1},0\leq n_{2}\leq N^{0}_{2},\\ [-\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}},\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}}]&\text{ otherwise}\end{cases}$ (21)
 where ${(e^{\bar{U}})}^{\mathrm{FFT}}$ is the numerical approximation of $e^{\bar{U}}$ calculated using discrete Fourier transforms, $C>0$, $\nu_{1},\nu_{2}>1$ and $\varepsilon_{n}>0$ for all $n$.
@@
 In this section, we justify our choices for the approximate objects constructed above by providing explicit formulas for the bounds of Theorem 3.1.
+Before stating the first bound, we recall that $N^{0}$ corresponds to the truncation dimension for sequences and $N$ for operators.
 Lemma 4.1 (Bound $\mathcal{Y}_{0}$).
@@
 $\|BF(\bar{U})\|_{2}^{2}=\|B^{N}F(\bar{U})\|_{2}^{2}+\|(\pi^{N^{0}}-\pi^{N})L\bar{U}+\pi_{N}G(\bar{U})\|_{2}^{2}.$ (24)
-Recall that $N^{0}$ corresponds to the truncation dimension for sequences and $N$ for operators.
-Note that the nonlinear part of $F(\bar{U})$ in $\|B^{N}F(\bar{U})\|_{2}^{2}$ can be calculated using (21).
+Note that the nonlinear part of $F(\bar{U})$ in $\|B^{N}F(\bar{U})\|_{2}^{2}$ can be calculated and bounded using (21).
 For the second term in the right-hand side of (24) we write
@@
 Note that for the second term in the $\mathcal{Y}_{0}$-bound, we can use (21) to state the interval that contains this term as follows:
-$\|(\pi^{N^{0}}-\pi^{N})(L\bar{U}+G(\bar{U}))\|_{2}^{2}\in\sum_{\begin{subarray}{c}n\in I_{N^{0}}\\ n\notin I_{N}\end{subarray}}\alpha_{n}\left(l(\tilde{n})(\bar{U})_{n}+(e^{\bar{U}})^{\mathrm{FFT}}+C[-\varepsilon_{n},\varepsilon_{n}]\right)^{2}.$
+$\|(\pi^{N^{0}}-\pi^{N})(L\bar{U}+G(\bar{U}))\|_{2}^{2}\in\sum_{\begin{subarray}{c}n\in I_{N^{0}}\\ n\notin I_{N}\end{subarray}}\alpha_{n}\left(l\left(\frac{n}{2d}\right)\bar{U}_{n}+(e^{\bar{U}})_{n}^{\mathrm{FFT}}+C[-\varepsilon_{n},\varepsilon_{n}]\right)^{2}.$
 The right-hand side is what is used in the (interval arithmetic) computation of the $\mathcal{Y}_{0}$-bound in [31].
@@
 The proof is given in [5]. Note that the difference in the formulas lies in the fact that $\sup_{\xi\in{\mathbb{R}}^{2}}\frac{1}{l(\xi)}=\frac{1}{1-\frac{c^{4}}{4}}$ in the case of (2). ∎
-First, we recall the computation of the bound $Z_{1}$, which is given in Lemma 3.9 from [7].
+We recall the computation of the bound $Z_{1}$, which is given in Lemma 3.9 from [5].
 Lemma 4.5.
@@
 Then, we obtain that
-$\mathcal{Z}_{u,1}$ $\leq 2d_{1}(\bar{V}^{N},\bar{V}^{N}*E)_{2},$
-$\mathcal{Z}_{u,2}$ $\leq 2d_{1}(\bar{V}^{N},V^{N}*E)_{2}+C(d_{1})e^{-2ad_{1}}(\bar{V}^{N},\bar{V}^{N}*E)_{2},$
+$\mathcal{Z}_{u,1}$ $\leq\bigg(2d_{1}(\bar{V}^{N},\bar{V}^{N}*E)_{2}\bigg)^{\frac{1}{2}},$
+$\mathcal{Z}_{u,2}$ $\leq\bigg(2d_{1}(\bar{V}^{N},V^{N}*E)_{2}+C(d_{1})e^{-2ad_{1}}(\bar{V}^{N},\bar{V}^{N}*E)_{2}\bigg)^{\frac{1}{2}},$
 and that $\mathcal{Z}_{u}=\max\left\{1,\|B^{N}\|_{2}\right\}\sqrt{\mathcal{Z}_{u,1}^{2}+\mathcal{Z}_{u,2}^{2}}$ satisfies (25).
@@
 $\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{n_{2}}^{-1}v_{n_{2}}\|_{2}^{2}\leq C_{n_{2}}^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\left(\int_{-d_{1}}^{d_{1}}e^{-a_{n_{2}}|x-y|}|v_{n_{2}}(x)|dx\right)^{2}dy$ (33)
-for all $n_{2}\in\mathbb{N}_{0}$, where we have used (28). Using that $v=\bar{v}^{N}u$, we have
+for all $n_{2}\in\mathbb{N}_{0}$, where we have used (28). Noting that $v=\bar{v}^{N}u$, we have
 $v_{n_{2}}=\sum_{k\in\mathbb{Z}}\bar{v}^{N}_{|n_{2}-k|}u_{|k|}.$
@@
 $\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{n_{2}}^{-1}v_{n_{2}}\|_{2}^{2}\leq C_{n_{2}}^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\left(\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}e^{-a_{n_{2}}|x-y|}|\bar{v}^{N}_{|n_{2}-k|}(x)u_{|k|}(x)|dx\right)^{2}dy.$
-Now, using the Cauchy-Schwarz inequality, we get
+Now, applying the Cauchy-Schwarz inequality, we get
 $\|\mathbb{1}_{\mathbb{R}\setminus(-d_{1},d_{1})}\mathbb{L}_{n_{2}}^{-1}v_{n_{2}}\|_{2}^{2}$
@@
 $\leq$~ $C_{n_{2}}^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}e^{-2a_{n_{2}}|x-y|}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}dxdy$ (34)
-since $\|u\|_{2}=1.$ Now, using Fubini’s theorem, we have
+since $\|u\|_{2}=1.$ Employing Fubini’s theorem, we have
 $\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\sum_{k\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}e^{-2a_{n_{2}}|x-y|}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}dxdy=\int_{-d_{1}}^{d_{1}}\sum_{k\in\mathbb{Z}}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\ e^{-2a_{n_{2}}|x-y|}dydx.$
-Using the proof of Theorem 3.9 from [6] again, we get
+Recalling the proof of Theorem 3.9 from [4], we get
 $\int_{{\mathbb{R}}\setminus(-d_{1},d_{1})}\ e^{-2a_{n_{2}}|x-y|}dy=\frac{e^{-2a_{n_{2}}d_{1}}\cosh(2a_{n_{2}}x)}{a_{n_{2}}}.$
@@
 $\tilde{E}_{n_{1}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{a_{n_{2}}(-1)^{n_{1}}(1-e^{-4a_{n_{2}}d_{1}})}{d_{1}(4a_{n_{2}}^{2}+(2\pi\tilde{n}_{1})^{2})}$
-for all $n_{1}\in\mathbb{N}_{0}$ and all $n_{2}\in\mathbb{N}_{0}$. Using Parseval’s identity, this implies that
+for all $n_{1}\in\mathbb{N}_{0}$ and all $n_{2}\in\mathbb{N}_{0}$. Applying Parseval’s identity, this implies that
 $\mathcal{Z}_{u,1}^{2}$ $\leq\sum_{k\in\mathbb{Z}}\sum_{n_{2}\in\mathbb{Z}}\int_{-d_{1}}^{d_{1}}|\bar{v}^{N}_{|n_{2}-k|}(x)|^{2}\frac{C_{|n_{2}|}^{2}e^{-2a_{n_{2}}d_{1}}\cosh(2a_{n_{2}}x)}{a_{n_{2}}}dx$
 $=2d_{1}(\bar{V}^{N},\bar{V}^{N}*E)_{2}.$ (35)
-This concludes the proof for $\mathcal{Z}_{u,1}$.
-For the computation of $\mathcal{Z}_{u,2}$ we use the formulas provided in [6]. Indeed, using the proof of Theorem 3.9, we can define $g_{n}$ as
+This concludes the inequality for $\mathcal{Z}_{u,1}$.
+For the computation of $\mathcal{Z}_{u,2}$ we use the formulas provided in [4]. Indeed, applying the proof of Theorem 3.9 again, we can define $g_{n}$ as
 $g_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2d_{1}}\int_{-d_{1}}^{d_{1}}\mathbb{L}_{n_{2}}v_{n_{2}}(x)e^{-2\pi i\tilde{n}_{1}x}dx$
-for all $n\in\mathbb{Z}^{2}$, where $v_{n_{2}}=v_{-n_{2}}$ if $n_{2}<0$. Using the proof of Theorem 3.9 in [6], we have that
+for all $n\in\mathbb{Z}^{2}$, where $v_{n_{2}}=v_{-n_{2}}$ if $n_{2}<0$. Utilizing the proof of Theorem 3.9 in [4], we have that
 $\mathcal{Z}_{u,2}^{2}=2d_{1}\sum_{n\in\mathbb{Z}^{2}}|g_{n}|^{2}=\int_{{\mathbb{R}}\setminus((-d_{1},d_{1})\cup(-d_{1},d_{1})+2d_{1}n_{1})}\mathbb{L}_{n_{2}}v_{n_{2}}(x)\mathbb{L}_{n_{2}}v_{n_{2}}(x-2d_{1}n_{1})dx.$
@@
 $\sum_{n_{2}\in\mathbb{Z}}\int_{{\mathbb{R}}\setminus((-d_{1},d_{1})\cup(-d_{1},d_{1})+2d_{1}n_{1})}\mathbb{L}_{n_{2}}v_{n_{2}}(x)\mathbb{L}_{n_{2}}v_{n_{2}}(x-2d_{1}n_{1})dx=\sum_{n_{2}\in\mathbb{Z}}\int_{{\mathbb{R}}\setminus((-d_{1},d_{1})}\left|\mathbb{L}_{n_{2}}v_{n_{2}}(x)\right|^{2}dx=\mathcal{Z}_{u,1}^{2}.$
-Using the proof of Lemma 6.5 in [6], we obtain that
+With the help of the proof of Lemma 6.5 in [4], we obtain that
 $\mathcal{Z}_{u,2}^{2}\leq\mathcal{Z}_{u,1}^{2}+\sum_{n\in\mathbb{Z}^{2},n_{1}\neq 0}C_{n_{2}}^{2}\int_{(-d_{1},d_{1})^{2}}|v_{n_{2}}(x)v_{n_{2}}(z)|I_{n}(x,z)dxdz$
@@
 5 Stability Analysis
-In this section, we aim at investigating the stability of traveling wave solutions to (2), of which the existence will be constructively established in Section 6.
+In this section, we aim at investigating the stability of traveling wave solutions to (2) with Neumann boundary conditions (3), which is the traveling wave formulation associated with (1).
+The existence of these traveling waves will be constructively established in Section 6.
 In particular, we assume that we know the existence of a solution $\tilde{u}\in\mathcal{H}$ to (2) such that
@@
 Moreover, let us define the functional $E:X\to{\mathbb{R}}$ as
-$E\begin{pmatrix}u\\ v\end{pmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\Omega}\frac{1}{2}v(x)^{2}+\frac{1}{2}(\Delta u(x))^{2}+e^{u(x)}-u(x)-1dx$ (37)
+$E\begin{pmatrix}u\\ v\end{pmatrix}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\int_{\Omega}\left[\frac{1}{2}v(x)^{2}+\frac{1}{2}(\Delta u(x))^{2}+e^{u(x)}-u(x)-1\right]dx$ (37)
 Then, (1) can be written in the form
@@
 Let $\tilde{u}$ be a traveling wave solution to (2) in $\mathcal{H}$ and let $\tilde{U}$ be defined as in (39).
-Then $\tilde{u}$ is said to be orbitally stable if the solution $U$ to (38) with initial condition $U(0)=U_{0}$ exists for all positive times and if for all $\epsilon>0$, there exists $\delta>0$ such that if $\|U_{0}-\tilde{U}\|_{X}<\delta$, then $\inf_{s\in{\mathbb{R}}}\|U(t)-T(s)\tilde{U}\|_{X}<\epsilon.$
-$\tilde{u}$ is called orbitally unstable if it is not orbitally stable.
+Then $\tilde{u}$ is said to be orbitally stable if the solution $U$ to (38) with initial condition $U(0)=U_{0}$ exists for all positive times and if for all $\epsilon>0$, there exists $\delta>0$ such that if $\|U_{0}-\tilde{U}\|_{X}<\delta$, then $\inf_{s\in{\mathbb{R}}}\|U(t)-T(s)\tilde{U}\|_{X}<\epsilon.$ Furthermore, $\tilde{u}$ is called orbitally unstable if it is not orbitally stable.
 […]
@@
 Remark 5.2.
-Using the notations of [18], $\tilde{w}$ in Lemma 5.1 corresponds to $v_{c}$ in Section 4.3 of [18].
-Moreover, $\theta$ corresponds to $d^{\prime\prime}(c)$, after using integrations by parts (cf.
-Section 4.3 of [18]).
+A solution is not only valid for a fixed parameter $c$, but also persists under variations of $c$.
+Differentiating with respect to $c$ yields $\tilde{w}$ in Lemma 5.1, which is called $v_{c}$ in [18].
+To determine (in)stability, the authors of [18] introduce a functional $d(c)$ and compute its second derivative $d^{\prime\prime}(c)$ in Section 4.3 of their paper, which corresponds to $\theta$ in our Lemma 5.1 after applying integrations by parts.
 […]
@@
 $\leq e^{\kappa_{2}r_{0}}\|e^{\bar{U}}\|_{1}\|u\|_{2}.$
-This concludes the proof. ∎
+This concludes the proof after a short algebraic manipulation. ∎
 […]
@@
 We define $S$ as the diagonal part of the matrix $\mathcal{D}$ and $R=\mathcal{D}-S$.
-Moreover, we define $\lambda_{k}$ as the diagonal entries of $S$, that is
+Moreover, we define $\lambda_{k}$ as the diagonal entries of $S$, that is,
 $(SU)_{k}=\lambda_{k}U_{k}\text{ for all }k\in\mathbb{N}_{0}^{2}.$ (46)
@@
 Lemma 5.4.
-Let $n_{2}\in\mathbb{N}_{0}$, and let us define $a_{n_{2}}(\mathcal{J})$, $b_{n_{2}}(\mathcal{J})$ and ${C}_{n_{2}}(\mathcal{J})$ as
+Let $n_{2}\in\mathbb{N}_{0}$, and let us define $a_{n_{2}}(\mathcal{J})$, $b_{n_{2}}(\mathcal{J})$ and ${C}_{n_{2}}(\mathcal{J})$ as (cf. Lemma 4.6)
 $a_{n_{2}}(\mathcal{J})$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\sqrt{4(1+\tilde{n}_{2}^{2}c^{2}-\delta_{0})-c^{4}}}{2\sqrt{c^{2}-2\tilde{n}_{2}^{2}+\sqrt{(c^{2}-2\tilde{n}_{2}^{2})^{2}+4(1+c^{2}\tilde{n}_{2}^{2}-\delta_{0})-c^{4}}}},$ (47)
@@
 $C_{n_{2}}(\mathcal{J})$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2(a_{n_{2}}(\mathcal{J})^{2}+b_{n_{2}}(\mathcal{J})^{2})^{\frac{1}{2}}\pi\sqrt{4(1+\tilde{n}_{2}^{2}c^{2}-\delta_{0})-c^{4}}}.$ (49)
-Moreover, given $\mu\in\mathcal{J}$, we define $f_{n_{2},\mu}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{F}^{-1}(\frac{1}{l_{n_{2}}-\mu})$.
-Then, we obtain that
+Given $\mu\in\mathcal{J}$, we define $f_{n_{2},\mu}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathcal{F}^{-1}(\frac{1}{l_{n_{2}}-\mu})$. Then, we obtain that
 $\sup_{\mu\in\mathcal{J}}|f_{n_{2},\mu}(x)|\leq C_{n_{2}}(\mathcal{J})e^{-a_{n_{2}}(\mathcal{J})|x|},~~\text{ for all }x\in{\mathbb{R}}.$ (50)
@@
 $C_{n_{2}}(\mu)$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{1}{2(a_{n_{2}}(\mu)^{2}+b_{n_{2}}(\mu)^{2})^{\frac{1}{2}}\pi\sqrt{4(1+\tilde{n}_{2}^{2}c^{2}-\mu)-c^{4}}}.$
-Then, one can prove that both $a_{n_{2}}$ and $b_{n_{2}}$ are decreasing functions of $\mu$, leading to the desired result. ∎
+Subsequently, one can prove that both $a_{n_{2}}$ and $b_{n_{2}}$ are decreasing functions of $\mu$, leading to the desired result. ∎
 […]
 Lemma 5.5.
-Let $t>-\lambda_{min}$ and let $\mathcal{J}=[\lambda_{min},\delta_{0}]$. Moreover, let us define $E=(E_{n})_{n\in\mathbb{N}_{0}^{2}}$ as follows:
+Let $t>-\lambda_{min}$ and let $\mathcal{J}=[\lambda_{min},\delta_{0}]$. We define $E=(E_{n})_{n\in\mathbb{N}_{0}^{2}}$ as
 $E_{n}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{(C_{n_{2}}(\mathcal{J}))^{2}a_{n_{2}}(\mathcal{J})(-1)^{n_{1}}(1-e^{-4a_{n_{2}}(\mathcal{J})d_{1}})}{d_{1}(4a_{n_{2}}(\mathcal{J})^{2}+(2\pi\tilde{n}_{1})^{2})}.$ (51)
@@
 $a(\mathcal{J})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\inf_{n_{2}\in\mathbb{N}_{0}}a_{n_{2}}(\mathcal{J})~\text{ and }~C(d_{1})\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}4d_{1}+\frac{4e^{-a(\mathcal{J})d_{1}}}{a(\mathcal{J})(1-e^{-\frac{3}{2}a(\mathcal{J})d_{1}})}+\frac{2}{a(\mathcal{J})(1-e^{-2a(\mathcal{J})d_{1}})}.$
-Now, let us introduce various constants satisfying the following
+We introduce various constants satisfying the following inequalities:
 $\mathcal{Z}_{u,1}$ $\geq 4d_{1}(\bar{V},\bar{V}*E),$
@@
 $Z_{1,3}$ $\geq\|\pi^{N}(S+tI)^{-1}R\pi^{N}\|_{2},~~~~Z_{1,4}\geq\|\pi^{N}(S+tI)^{-1}R\pi_{N}\|_{2}.$
-If
+Assuming
 $\mathcal{C}_{1}r_{0}<1,$ (52)
-we define $\beta_{1}$ as $\beta_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{u,1}+\mathcal{C}_{1}r_{0}}{1-\mathcal{C}_{1}r_{0}}.$ If in addition
-$1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}>0,$ (53)
-then we define
+we define $\beta_{1}$ as $\beta_{1}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{\mathcal{Z}_{u,1}+\mathcal{C}_{1}r_{0}}{1-\mathcal{C}_{1}r_{0}}.$ Furthermore, we assume that
+$1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}>0.$ (53)
+Then, we define
 $\beta_{2}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\frac{Z_{1,1}+(\mathcal{Z}_{u,2}+(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0})\|P^{N}\|_{2}}{1-Z_{1,2}-\mathcal{Z}_{u,2}-(1+\beta_{1}^{2})^{\frac{1}{2}}\mathcal{C}_{1}r_{0}},$
@@
 $\epsilon_{n}$ $\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\max\{\epsilon_{n}^{(q)},\epsilon_{n}^{(\infty)}\}$
+for all $n\in\mathbb{N}_{0}^{2}$. Let $k\in\mathbb{N}$ and $I\subset\mathbb{N}_{0}^{2}$ such that $|I|=k$.
 […]
 Proof.
-The proof of the lemma follows the ones of Lemmas 4.1 and 4.2 in [4].
-Now, we comment on the computation of the bounds.
-Following Lemma 4.1 in [4], we want to compute $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2}$ such that
+The proof of the lemma follows the ones of Lemmas 4.1 and 4.2 in [7]. Following Lemma 4.1 in [7], we want to compute $\mathcal{Z}_{u,1},\mathcal{Z}_{u,2}$ such that
 $\mathcal{Z}_{u,1}$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\Omega\setminus\Omega_{0}}\left(\mathbb{L}-\mu I\right)^{-1}D\mathbb{G}(\bar{u})\right\|_{2},$
 $\mathcal{Z}_{u,2}$ $\geq\sup_{\mu\in\mathcal{J}}\|\mathbb{1}_{\Omega_{0}}\left(\Gamma^{\dagger}\left((L-\mu I)^{-1}\right)-(\mathbb{L}-\mu I)^{-1}\right)D\mathbb{G}(\bar{u})\|_{2}.$
-Now, noticing that $\left(\mathbb{L}_{n_{2}}-\mu I\right)^{-1}$ is the convolution operator associated to the function $f_{n_{2},\mu}$ defined in Lemma 5.4, we combine Lemma 4.7 and Lemma 5.4 to obtain that
-$2d_{1}(\bar{V},\bar{V}*E)$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\Omega\setminus\Omega_{0}}\left(\mathbb{L}-\mu I\right)^{-1}D\mathbb{G}(\bar{u})\right\|_{2}$
+Noticing that $\left(\mathbb{L}_{n_{2}}-\mu I\right)^{-1}$ is the convolution operator associated to the function $f_{n_{2},\mu}$ defined in Lemma 5.4, we combine Lemma 4.7 and Lemma 5.4 to obtain that
+$2d_{1}(\bar{V},\bar{V}*E)$ $\geq\sup_{\mu\in\mathcal{J}}\left\|\mathbb{1}_{\Omega\setminus\Omega_{0}}\left(\mathbb{L}-\mu I\right)^{-1}D\mathbb{G}(\bar{u})\right\|_{2},$
 $2d_{1}(\bar{V},\bar{V}*E)_{2}+C(d_{1})e^{-2ad_{1}}(\bar{V},\bar{V}*E)_{2}$ $\geq\sup_{\mu\in\mathcal{J}}\|\mathbb{1}_{\Omega_{0}}\left(\Gamma^{\dagger}\left((L-\mu I)^{-1}\right)-(\mathbb{L}-\mu I)^{-1}\right)D\mathbb{G}(\bar{u})\|_{2}.$
@@
 $\|\pi^{N}(S+tI)^{-1}P^{-1}(L-\mu I)\|_{2}\leq\|\pi^{N}(S+tI)^{-1}P^{-1}(L-\delta_{0}I)\|_{2}\|(L-\delta_{0}I)^{-1}(L-\mu I)\|_{2}.$
-But then, we have
+Then, we have
 $\|(L-\delta_{0}I)^{-1}(L-\mu I)\|_{2}=\sup_{n\in\mathbb{N}_{0}^{2}}\frac{l(\tilde{n})-\mu}{l(\tilde{n})-\delta_{0}}=\sup_{n\in\mathbb{N}_{0}^{2}}1+\frac{\delta_{0}-\mu}{l(\tilde{n})-\delta_{0}}=\frac{1-\frac{c^{4}}{4}-\mu}{1-\frac{c^{4}}{4}-\delta_{0}}$
@@
 Then, recalling $C_{n_{2}}(\mu)$ from the proof of Lemma 5.4, one can prove that $\mu\mapsto C_{n_{2}}(\mu)\frac{1-\frac{c^{4}}{4}-\mu}{1-\frac{c^{4}}{4}-\delta_{0}}$ is increasing in $\mu$ and consequently the maximum is reached at $\mu=\delta_{0}$ when considering $\mu\in\mathcal{J}$.
-This concludes the obtained inequality for $\mathcal{Z}_{u,3}$.
-Now, the computations of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ follow directly from the proof of Lemma 4.3.
+This concludes the inequality for $\mathcal{Z}_{u,3}$.
+The computations of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ follow directly from the proof of Lemma 4.3.
 In particular, we also use that $\|(\mathbb{L}-\mu I)^{-1}\|_{2}\leq\frac{1}{1-\frac{c^{4}}{4}-\delta_{0}}$ for all $\mu\in\mathcal{J}$.
@@
 Let $r_{0}>0$ and let $\tilde{u}\in\mathcal{H}$ be a solution to (2) such that $\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}$, where $\bar{u}\in\mathcal{H}$ is constructed as in (20).
-Let $\overline{W}\in\pi^{N}\ell^{2}$ such that $\overline{w}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(\overline{W})\in\mathcal{H}$.
+Choose $\overline{W}\in\pi^{N}\ell^{2}$ such that $\overline{w}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\gamma^{\dagger}(\overline{W})\in\mathcal{H}$.
 Let $\epsilon_{0},\epsilon>0$ be defined as
@@
 $\theta=\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}\bar{u}+\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})+\int_{\Omega}(\tilde{u}-\overline{u}+2c\tilde{w}-2c\overline{w})\partial_{x_{1}}^{2}\tilde{u}$,
-where $\overline{w}\in\mathcal{H}$ is an approximation of $\tilde{w}$. In particular, using Cauchy-Schwarz inequality, we have
+where $\overline{w}\in\mathcal{H}$ is an approximation of $\tilde{w}$. In particular, employing Cauchy-Schwarz inequality, we have for the second term
 $\left|\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})\right|$ $\leq\|\bar{u}+2c\overline{w}\|_{2}\|\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})\|_{2}$
@@
 $\leq\|\bar{u}+2c\overline{w}\|_{2}\|\partial_{x_{1}}^{2}\mathbb{L}^{-1}\|_{2}r_{0}$
-using that $\|\mathbb{L}(\tilde{u}-\bar{u})\|_{2}=\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}.$ Moreover, using the Fourier transform, we have that
+using that $\|\mathbb{L}(\tilde{u}-\bar{u})\|_{2}=\|\tilde{u}-\bar{u}\|_{\mathcal{H}}\leq r_{0}.$ Applying the Fourier transform, we have for the second norm on the right-hand side that
 $\|\partial_{x_{1}}^{2}\mathbb{L}^{-1}\|_{2}\leq\sup_{\xi\in{\mathbb{R}}^{2}}\frac{(2\pi\xi_{1})^{2}}{l(\xi)}\leq\frac{1}{2-c^{2}}.$
-On the other hand, we have
+Furthermore, we have for the third term of $\theta$ that
 […]
 $\leq(\|\tilde{u}-\overline{u}\|_{2}+2c\|\tilde{w}-\overline{w}\|_{2})(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\|\partial_{x_{1}}^{2}(\tilde{u}-\bar{u})\|_{2})$
-$\leq(\kappa_{1}r_{0}+2c\|\tilde{w}-\overline{w}\|_{2})(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\frac{r_{0}}{2-c^{2}}).$
-Consequently, it remains to compute $\|\tilde{w}-\overline{w}\|_{2}$. By definition of $\tilde{w}$ in (40), we have
-$\|\tilde{w}-\overline{w}\|_{2}=\|-2cD\mathbb{F}(\tilde{u})\partial_{x_{1}}^{2}\tilde{u}-\overline{w}\|_{2}\leq 2c\|D\mathbb{F}(\tilde{u})^{-1}\|_{2}\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
-Now, we have
-$\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}\leq\|\partial_{x_{1}}\bar{u}+\frac{1}{2c}D\mathbb{F}(\bar{u})\overline{w}\|_{2}+\|\partial_{x_{1}}^{2}(\bar{u}-\tilde{u})\|_{2}+\frac{1}{2c}\|D\mathbb{F}(\bar{u})\overline{w}-D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
-Moreover, we obtain that
+$\leq\left(\kappa_{1}r_{0}+2c\|\tilde{w}-\overline{w}\|_{2}\right)\left(\|\partial_{x_{1}}^{2}\bar{u}\|_{2}+\tfrac{r_{0}}{2-c^{2}}\right).$
+It remains to estimate $\|\tilde{w}-\overline{w}\|_{2}$. By definition of $\tilde{w}$ in (40), we have
+$\|\tilde{w}-\overline{w}\|_{2}=\|-2cD\mathbb{F}(\tilde{u})\partial_{x_{1}}^{2}\tilde{u}-\overline{w}\|_{2}\leq 2c\|D\mathbb{F}(\tilde{u})^{-1}\|_{2}\|\partial_{x_{1}}\tilde{u}+\tfrac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
+For the second norm on the right-hand side
+$\|\partial_{x_{1}}\tilde{u}+\frac{1}{2c}D\mathbb{F}(\tilde{u})\overline{w}\|_{2}\leq\|\partial_{x_{1}}\bar{u}+\frac{1}{2c}D\mathbb{F}(\bar{u})\overline{w}\|_{2}+\left|\partial_{x_{1}}^{2}(\bar{u}-\tilde{u})\right\|_{2}+\frac{1}{2c}\|D\mathbb{F}(\bar{u})\overline{w}-D\mathbb{F}(\tilde{u})\overline{w}\|_{2}.$
+For the last term, we estimate
 $\|D\mathbb{F}(\bar{u})\overline{w}-D\mathbb{F}(\tilde{u})\overline{w}\|_{2}=\|(e^{\bar{u}}-e^{\tilde{u}})\overline{w}\|_{2}\leq\|1-e^{\tilde{u}-\bar{u}}\|_{2}\|e^{\bar{u}}\overline{w}\|_{\infty}$
 $\leq(1-e^{\kappa_{2}r_{0}})\|e^{\bar{U}}\overline{W}\|_{1}.$
-Finally, using that $\bar{u}=\gamma^{\dagger}(\bar{U})$ and $\overline{w}=\gamma^{\dagger}(\overline{W})$, we use Parseval’s identity to get
+Finally, using $\bar{u}=\gamma^{\dagger}(\bar{U})$ and $\overline{w}=\gamma^{\dagger}(\overline{W})$, we apply Parseval’s identity to get
 $\int_{\Omega}(\bar{u}+2c\overline{w})\partial_{x_{1}}^{2}\bar{u}=|{\Omega_{0}}|\left((\bar{U}+2c\overline{W}),\partial_{x_{1}}^{2}\bar{U}\right)_{2},$
-where $(\partial_{x_{1}}^{2}\bar{U})_{n}=-\frac{n_{1}^{2}\pi^{2}}{d_{1}^{2}}\bar{U}_{n}$ for all $(n_{1},n_{2})\in\mathbb{N}_{0}^{2}$. This concludes the proof. ∎
+where $(\partial_{x_{1}}^{2}\bar{U})_{n}=-\frac{n_{1}^{2}\pi^{2}}{d_{1}^{2}}\bar{U}_{n}$ for all $(n_{1},n_{2})\in\mathbb{N}_{0}^{2}$. After algebraic manipulation, this concludes the proof. ∎
 Using the previous result, we obtain a computer-assisted strategy for the enclosure of the value of $\theta$.
@@
 6 Results
-In this section, we present the main results of the paper, which consist of rigorous, computer-assisted bounds on numerical approximations to (2).
-These approximate solutions were obtained using Newton’s method, initialized with carefully constructed guesses based on approximations in the periodic setting described in [32].
+In this section, we present computer-assisted proofs for the existence of solutions to (2) with Neumann boundary conditions (3).
+To do so, we begin with approximate solutions obtained by Newton’s method and initialized with carefully constructed guesses based on approximations in the periodic setting described in [32].
 To extend the initial guesses in the periodic case to the infinite setting, we project them into the kernel of matrix $\mathcal{T}$, as described in Section 3.1.
@@
 Finitely many cosine coefficients of the approximate solution $\bar{u}$, visualized in Figure 1, can be found in the code in [31].
-We report the values for the bounds in floating point numbers displaying four decimals.
+We report the values for the bounds in floating point numbers displaying finitely many decimals.
 We have
@@
 The stability part follows similarly as the one presented in the proof of Theorem 6.1.
-In this case, we provide the existence of exactly two negative eigenvalues by proving that two Gershgorin intervals of Lemma 5.5 are in the negative part of $\mathbb{R}$}, disjoint from on another, and disjoint from the rest of the intervals.
+In this case, we provide the existence of exactly two negative eigenvalues by proving that two Gershgorin intervals of Lemma 5.5 are in the negative part of $\mathbb{R}$}, disjoint from one another, and disjoint from the rest of the intervals.
 ∎
+The next theorem and its proof are very similar to the previous ones. However, the stability analysis is more involved, so we choose to present the full proof again.
 Theorem 6.3.
@@
 This implies that we can improve our choice for $\lambda_{min}$ and choose $\lambda_{min}=-0.45$.
-Choosing $t=0.5$ in Lemma 5.3 and computing the bounds again, we get tighter radiae for the Gershgoring intervals and obtain disjoint intervals on the negative part of $\mathbb{R}$}.
+Choosing $t=0.5$ in Lemma 5.3 and computing the bounds again, we get tighter radii for the Gershgorin intervals and obtain disjoint intervals on the negative part of $\mathbb{R}$}.
 The stability part then follows similarly as the one presented in the proof of Theorem 6.1.
@@
 The increased computation time associated with our first theorem arises from selecting a larger value for $N_{0}$, which in turn yields a tighter error bound.
+7 Acknowledgements
+Matthieu Cadiot was supported by the ANR project CAPPS: ANR-23-CE40-0004-01 and by the FMJH : ANR-22-EXES-0013. Lindsey van der Aalst was partially supported by NWO grant 613.009.132.
 8 Appendix
-In the next lemma we provide some technical results about $l_{n}$.
+In the next lemma we provide some technical results about $l_{n}$, where we recall from (6) that
+$l_{n_{2}}(\xi_{1})=(2\pi)^{4}(\xi_{1}^{2}+\tilde{n}_{2}^{2})^{2}-c^{2}(2\pi\xi_{1})^{2}+1$
+and that $\tilde{n}_{2}=\frac{n_{2}}{2d_{2}}$.
 Lemma 8.1.
@@
 Noticing that $2\pi\tilde{n}_{2}\geq c$ is equivalent to $n_{2}\geq\frac{d_{2}c}{\pi}$, we have
-$2\sum_{n_{2}=N+1}^{\infty}\left\|\frac{1}{l_{n_{2}}}\right\|_{2}^{2}\leq\frac{(2d_{2})^{7}}{6N^{6}}\frac{5}{8(2\pi)^{7}}$
-where $N=\lceil{\frac{d_{2}c}{\pi}}\rceil.$
+$\sum_{n_{2}=N+1}^{\infty}\left\|\frac{1}{l_{n_{2}}}\right\|_{2}^{2}\leq\frac{(2d_{2})^{7}}{6N^{6}}\frac{5}{8(2\pi)^{7}}$
+where $N\in\mathbb{N}$ such that $N>\frac{d_{2}c}{\pi}.$